### v2

# Packing subgraphs in regular graphs

Abstract
An $H$-packing in a graph $G$ is a collection of pairwise vertex-disjoint copies of $H$ in $G$. We prove that for every $c>0$ and every bipartite graph $H$, any $\left\lfloor cn\right\rfloor$-regular graph $G$ admits an $H$-packing that covers all but a constant number of vertices. This resolves a problem posed by Kühn and Osthus in 2005. Moreover, our result is essentially tight: the conclusion fails if $G$ is not both regular and sufficiently dense, it is in general not possible to guarantee covering all vertices of $G$ by an $H$-packing, and if $H$ is non-bipartite then $G$ need not contain any copies of $H$.
We also prove that for all $c>0$, integers $t\geq 2$, and sufficiently large $n$, all the vertices of every $\lfloor cn\rfloor$-regular graph can be covered by vertex-disjoint subdivisions of $K_{t}$. This resolves another problem of Kühn and Osthus from 2005, which goes back to a conjecture of Verstraëte from 2002.
Our proofs combine novel methods for balancing expanders and super-regular subgraphs with a number of powerful techniques including properties of robust expanders, regularity lemma, and blow-up lemma.
1 Introduction
Given two graphs $F$ and $G$, an $F$-packing (or an $F$-tiling) in $G$ is a collection of pairwise vertex-disjoint copies of $F$ in $G$. An $F$-packing in $G$ is called perfect if it covers all the vertices of $G$. Note that when $F$ consists of a single edge, an $F$-packing is a graph matching. Tutte’s theorem characterizes those graphs which have a perfect $F$-packing if $F$ is an edge, but no such characterisation is known for other connected graphs $F$. In fact, it is known [19] that the decision problem of whether a graph $G$ has a perfect $F$-packing is NP-complete if and only if $F$ has a component which contains at least three vertices. So it is natural to seek simple sufficient conditions that guarantee the existence of a perfect $F$-packing in a given graph.
A fundamental result in this area is the Hajnal–Szemerédi theorem [13] from 1970 which states that every graph whose order $n$ is divisible by $t$ and whose minimum degree is at least $(1-\frac{1}{t})n$ contains a perfect $K_{t}$-packing. For arbitrary graphs, Komlós, Sárközy and Szemerédi [22] showed that for every graph $F$, there exists a constant $C=C(F)$ such that every graph $G$ whose order $n$ is divisible by $|V(F)|$ and whose minimum degree is at least $(1-\frac{1}{\chi(F)})n+C$ contains a perfect $F$-packing. It turns out that there are graphs $F$ for which the bound on the minimum degree can be improved significantly i.e., one can often replace $\chi(F)$ by a smaller parameter called the critical chromatic number $\chi_{cr}(F)$ of $F$, which is defined as $\chi_{cr}(F)\coloneqq(\chi(F)-1)\frac{|F|}{|F|-\sigma(F)}$. Here $\sigma(F)$ denotes the minimum size of the smallest colour class in an optimal colouring of $F$. It is easy to see that $\chi(F)-1<\chi_{cr}(F)\leq\chi(F)$. In 2000, Komlós [20] showed that the critical chromatic number is the parameter that governs the existence of almost perfect packings in graphs of large minimum degree. Finally, in 2009, Kühn and Osthus [32] determined, up to an additive constant, the minimum degree of a graph $G$ that ensures the existence of a perfect $F$-packing in $G$, for every graph $F$. More precisely, they proved that for every graph $F$, either its critical chromatic number or its chromatic number is the parameter which governs the existence of perfect $F$-packings in graphs of large minimum degree (where the exact classification depends on a parameter called the highest common factor of $F$).
1.1 $F$-packings in dense regular graphs
In view of the above results, rather surprisingly, Kühn and Osthus [31] showed that if we restrict our attention to packings in regular graphs, then any linear bound on the degree guarantees an almost perfect $F$-packing.
Theorem 1.1 (Kühn and Osthus [31]).
For every bipartite graph $F$ and every $0<c,\alpha\leq 1$, there exists $n_{0}=n_{0}(F,c,\alpha)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$ and $n\geq n_{0}$, has an $F$-packing that covers all but at most $\alpha n$ vertices of $G$.
It is easy to see that the restriction to (sufficiently) regular graphs $G$ in Theorem 1.1 is necessary. Indeed, if $G$ is, say, a complete bipartite graph $K_{n/4,3n/4}$ and $F$ is an edge, then it is clearly impossible to find an $F$-packing covering almost all the vertices of $G$. The restriction to bipartite graphs $F$ in Theorem 1.1 is also necessary because if $c\leq 1/2$, then $G$ could be a bipartite graph (in which case it cannot contain any subgraph $F$ which is non-bipartite).
This raises the natural question of whether the bound $\alpha n$ on the number of uncovered vertices in Theorem 1.1 can be lowered significantly to obtain an $F$-packing which is close to being perfect. Indeed, Kühn and Osthus [31] proposed the following general problem in 2005.
Problem 1.2 (Kühn and Osthus [31]).
Is it true that, for every bipartite graph $F$ and every $0<c\leq 1$, there is a constant $C=C(F,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has an $F$-packing that covers all but at most $C$ vertices of $G$?
Kühn and Osthus [31] resolved Problem 1.2 in the special case when the parts of $F$ have unequal sizes. More precisely, they showed the following.
Theorem 1.3 (Kühn and Osthus [31]).
For every bipartite graph $F$ whose vertex classes have different size and for every $0<c\leq 1$, there is a constant $C=C(F,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has an $F$-packing which covers all but at most $C$ vertices of $G$.
Despite extensive research on graph packings, this problem (Problem 1.2) remained open for nearly twenty years. Our main result is a complete resolution of Problem 1.2 as follows.
Theorem 1.4.
For every bipartite graph $F$ and every $0<c\leq 1$, there is a constant $C=C(F,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has an $F$-packing that covers all but at most $C$ vertices of $G$.
As noted in [31], it is necessary to allow for uncovered vertices in Theorem 1.4, and the upper bound $C$ on the number of uncovered vertices must depend on $c$ and $F$ (even if $n$ is divisible by $|V(F)|$), in sharp contrast to the result of Komlós, Sárközy and Szemerédi [22] mentioned earlier. Indeed, for instance, if $G$ is the vertex-disjoint union of cliques of order k|V(F)|1=d+1k|V(F)|-1=d+1 (for some positive integer $k$), then $G$ is $d$-regular, and it is easy to see that we must have at least $(|V(F)|-1)\cdot\frac{n}{d+1}$ uncovered vertices in any $F$-packing of $G$. Moreover, this example also shows that $d$ must be linear in $n$ in order for the number of uncovered vertices to be bounded by a constant $C$, so the requirement that the vertex degrees of $G$ are linear in $n$ in Theorem 1.4 is also necessary.
It is easy to reduce Theorem 1.4 to the special case when $F$ is a complete bipartite graph $K_{t,t}$. This follows, for example, from the fact that any bipartite graph $F$ can be perfectly packed into the complete bipartite graph $K_{|F|,|F|}$. To prove Theorem 1.4 when $F=K_{t,t}$, we start by partitioning the vertex set of our dense regular graph $G$ into a small number of expanders using a structural decomposition result from the work of Gruslys and Letzter [11], which goes back to the work of Kühn, Lo, Osthus and Staden [27, 28] on Hamilton cycles in dense regular graphs. These expanders are, in turn, closely related to a powerful notion called robust expanders that was introduced by Kühn, Osthus and Treglown [34] and has since been used to prove several longstanding conjectures (see, e.g., [30, 29, 33]).
One of the key contributions of our paper is a novel method for ‘balancing’ these expanders. By this, we mean transforming expanders that are bipartite (or close to being bipartite) with imbalanced part sizes into balanced bipartite expanders. This method thus enables us to obtain expanders that are either balanced bipartite or far from bipartite by removing a carefully constructed, small $K_{t,t}$-packing. Our main idea is to track a certain subgraph, which initially consists of the edges between expanders. We then sequentially move vertices between expanders and update the subgraph accordingly, ensuring that by the end, it has bounded maximum degree while still containing enough edges to admit a $K_{t,t}$-packing that balances the expanders (up to an additive constant, which is sufficient for our purposes). Note that while balancing expanders also plays a key role in [27, 28]—in resolving a problem of Bollobás and Häggkvist on Hamilton cycles in regular graphs—and in [11], which proves a conjecture of Magnant and Martin [39] for dense graphs, our approach differs significantly from the methods used in these works, so it is of independent interest. See Section 2.1 for further details on our balancing procedure.
After balancing the expanders, the original problem reduces to finding an almost perfect $K_{t,t}$-packing either in a balanced bipartite expander or in an expander that is far from bipartite. To address the latter case, our first goal is to apply the regularity lemma to obtain vertex-disjoint super-regular pairs covering all but a small set of exceptional vertices. For this, rather than directly seeking a perfect matching in the reduced graph, we exploit the Hamiltonicity of expanders to find a perfect fractional matching (inspired by [17, 16, 25]). This fractional matching is used to construct a collection of edges and odd cycles covering all vertices of the reduced graph. We then refine the clusters by splitting them each into two equal parts and then remove a small number of vertices from the new clusters to obtain vertex-disjoint super-regular pairs covering all but a small exceptional set of vertices, as required.
Second, we cover the exceptional vertices with a $K_{t,t}$-packing, while also ensuring that none of the super-regular pairs are overused. Third, we construct another small $K_{t,t}$-packing, whose purpose is to ensure that the number of remaining vertices in each part of every super-regular pair is divisible by $t$. Ideally, at this point, one could attempt to complete the proof by applying the blow-up lemma to find a perfect $K_{t,t}$-packing covering the uncovered portion of each super-regular pair. However, this is not immediately feasible because the number of remaining vertices in the two parts of a super-regular pair may differ. To address this, we introduce another novel balancing strategy: we construct a suitable matching within a small subgraph induced by the uncovered vertices, once again exploiting the Hamiltonicity of expanders. This matching serves as a ‘template’ for constructing a carefully designed balancing $K_{t,t}$-packing, whose removal allows us to find a perfect $K_{t,t}$-packing covering all remaining vertices using the blow-up lemma, as intended. We give a detailed sketch of this balancing strategy in Section 2.2, and a more detailed outline of our methods in Section 2.
1.2 Packing subdivisions in dense regular graphs
Given a graph $F$, a subdivision of $F$, denoted by $\operatorname{TF}$, is a graph obtained by replacing the edges of $F$ with pairwise internally vertex-disjoint paths between the corresponding ends, whose interiors avoid the vertices of $F$. In this case, we refer to the vertices of $F$ as the branch vertices of $\operatorname{TF}$. For instance, a subdivision $\operatorname{TK}_{t}$ of the complete graph of order $t$ consists of $t$ (branch) vertices $\{v_{1},\ldots,v_{t}\}$ and $\binom{t}{2}$ pairwise internally vertex-disjoint paths $P_{i,j}$, $1\leq i<j\leq t$, such that $P_{i,j}$ joins $v_{i}$ and $v_{j}$ and avoids all other vertices in $\{v_{1},\ldots,v_{t}\}$.
The notion of subdivisions has played an important role in topological graph theory since the seminal result of Kuratowski [35] from 1930 showing that a graph is planar if and only if it does not contain a subdivision of $K_{5}$ or $K_{3,3}$. One of the most classical results in this area is due to Mader [38], who showed that there is some $d=d(t)$ such that every graph with average degree at least $d$ contains a subdivision of the complete graph $K_{t}$. Mader [38], and independently Erdős and Hajnal [9] conjectured that $d(t)=O(t^{2})$. In the 1990s, Komlós and Szemerédi [23, 24], and independently, Bollobás and Thomason [4] confirmed this conjecture. Since then, various extensions and strengthenings of this result have been studied, see, e.g., [10, 37, 40].
Given graphs $F$ and $G$, a $\operatorname{TF}$-packing in $G$ is a collection of pairwise vertex-disjoint copies of subdivisions of $F$ in $G$ (which are not required to be isomorphic). In 2002, Verstraëte [44] made the following conjecture.
Conjecture 1.5 (Verstraëte [44]).
For every graph $F$ and every $\eta>0$, there exists $d_{0}=d_{0}(F,\eta)$ such that, for all integers $d\geq d_{0}$, every $d$-regular graph $G$ of order $n$ contains a $\operatorname{TF}$-packing which covers all but at most $\eta n$ vertices of $G$.
Quite a lot of research has established Conjecture 1.5 in several natural special cases. The conjecture follows from a result of Kelmans, Mubayi and Sudakov [18] when $F$ is a tree, and in 2003, Alon [1] proved it when $F$ is a cycle. In 2005, Kühn and Osthus [31] observed that when $G$ is dense (i.e., $d=\Omega(n)$) the conjecture follows from their results on packings in dense regular graphs. Making significant progress towards the general case, Letzter, Methuku and Sudakov [36] applied their work on nearly-Hamilton cycles in sublinear expanders to show that the conjecture holds for all graphs $F$ when $G$ has degree at least polylogarithmic in $n$. Finally, very recently, Montgomery, Petrova, Ranganathan and Tan [42] resolved the conjecture in full.
For subdivisions of complete graphs $K_{t}$, Kühn and Osthus [31] went even further and showed that one can actually guarantee perfect packings in dense regular graphs for $t=4$ and $t=5$. More precisely, they showed that for every $0<c\leq 1$, there is a positive number $n_{0}=n_{0}(c)$ such that every $d$-regular graph $G$ of order $n$, with $d\geq cn$ and $n\geq n_{0}$, has a perfect $\operatorname{TK}_{t}$-packing for $t=4$ and $t=5$. By a result of Gruslys and Letzter [11], it follows that this also holds for $t=2$ and $t=3$.
In 2005, Kühn and Osthus asked whether the same holds for $t\geq 6$.
Problem 1.6 (Kühn and Osthus [31]).
Given $t\geq 6$ and $0<c\leq 1$, does every $d$-regular graph $G$ of order $n$, where $d\geq cn$ and $n$ is sufficiently large, have a perfect $\operatorname{TK}_{t}$-packing?
Kühn and Osthus’s result [31] actually guarantees a perfect $\operatorname{TK}_{t}$-packing, with $t\in\{4,5\}$, even when $G$ is only assumed to be almost regular. In contrast, for $t=3$ or $t\geq 6$, the restriction to regular graphs $G$ in Problem 1.6 is necessary. Indeed, it is easy to see that the complete bipartite graph $K_{m,m+1}$ does not have a perfect $\operatorname{TK}_{t}$-packing if $t=3$ (note that in this case a $\operatorname{TK}_{3}$-packing corresponds to a cycle partition, i.e., a collection of pairwise vertex-disjoint cycles covering all vertices). Interestingly, Kühn and Osthus showed that $K_{m,m+1}$ also does not have a perfect $\operatorname{TK}_{t}$-packing if $t\geq 6$ (see Proposition 5.1 in [31]).
Our second result is the complete resolution of Problem 1.6. In fact, we prove a stronger result, showing that there is a perfect $\operatorname{TF}$-packing, for every graph $F$.
Theorem 1.7.
For every graph $F$ and $0<c\leq 1$, there is a positive number $n_{0}=n_{0}(c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$ and $n\geq n_{0}$, has a perfect $\operatorname{TF}$-packing.
It is unclear whether the requirement that the vertex degrees in $G$ are linear in $n$ in Theorem 1.7 is necessary. In this direction, Kühn and Osthus [31] gave an example showing that for all $t\geq 3$ the vertex degrees in $G$ must be at least $\sqrt{n/2}$ to guarantee a perfect $\operatorname{TK}_{t}$-packing in $G$, even if $G$ is regular. Interestingly, the case $t=2$ is different. It is a simple exercise to show that every $d$-regular graph admits a perfect $\operatorname{TK}_{2}$-packing, i.e., its vertex set can be partitioned into vertex-disjoint paths of positive length. (In fact, if $d$ is even, then by Petersen’s 2-factor theorem [43] it also admits a perfect $\operatorname{TK}_{3}$-packing.) It is therefore natural to ask how few paths are required to achieve such a partition. In this direction, Magnant and Martin [39] conjectured that the vertices of every $d$-regular graph on $n$ vertices can be covered by at most $n/(d+1)$ pairwise vertex-disjoint paths. This conjecture of Magnant and Martin was confirmed for $d\leq 5$ by Magnant and Martin themselves and for dense graphs $G$ by Gruslys and Letzter [11]. Montgomery, Müyesser, Pokrovskiy and Sudakov [41] showed that almost all vertices of every $n$-vertex $d$-regular graph can be covered by at most $n/(d+1)$ vertex-disjoint paths. Very recently, Christoph, Draganić, Girão, Hurley, Michel and Müyesser [6] proved that every $d$-regular graph can be covered by at most $2n/(d+1)$ vertex-disjoint paths, thereby proving the Magnant–Martin conjecture up to a constant factor and confirming a conjecture of Feige and Fuchs (see also [5]).
To prove Theorem 1.7, we begin by partitioning the vertex set of the dense regular graph $G$ into a small number of expanders, following the approach used in the proof of Theorem 1.4. In this setting, however, we employ a lemma from [11] that provides a small linear forest $H$ that balances the expanders, ensuring that each expander contains either zero or two leaves of $H$. We then modify this linear forest so that at most one component of the new linear forest is associated with each expander, with both leaves of the component contained within that expander. Next, we construct a pair of small, vertex-disjoint $F$-subdivisions in each expander, while also ensuring that their union remains balanced. Finally, we absorb the components of $H$, along with any remaining uncovered vertices in each expander, into these $F$-subdivisions. This absorption step again relies on the Hamiltonicity of expanders and a robust connectivity property, which guarantees that any two vertices can be joined by a short path that avoids a small set of forbidden vertices. See Section 7 for a more detailed sketch of these ideas.
1.3 Organisation of the paper
In the next subsection, we present the notation used throughout the paper. In Section 2, we outline a detailed proof sketch of our main result, Theorem 1.4, on almost perfect packings in dense regular graphs. In Section 3 we collect some tools and lemmas that are used throughout the paper. Key to our proof of Theorem 1.4 are two lemmas – Lemma 4.1 (for balancing expanders) and Lemma 5.1 (for packing $K_{t,t}$’s in expanders). We prove Lemma 4.1 in Section 4 and Lemma 5.1 in Section 5, and complete the proof of Theorem 1.4 in Section 6. In Section 7, we prove our second result, Theorem 1.7, which addresses perfect packings of subdivisions in dense regular graphs. A detailed sketch of its proof is given at the beginning of Section 7.
1.4 Notation
Let $[m]$ denote the set $\{1,\ldots,m\}$. We write $c=a\pm b$ if $a-b\leq c\leq a+b$. We use the $\ll$’ notation to state many of our results. We write $a\ll b$ to mean that there exists a non-decreasing function $f:(0,1]\mapsto(0,1]$ such that the subsequent result holds for all $a,b\in(0,1]$ with $a\leq f(b)$. We will not calculate these functions explicitly. Hierarchies with more constants are defined in a similar way and are to be read from right to left.
For a graph $G$, let $e(G)$ denote the number of edges in $G$ and, for two sets $X,Y\subseteq V(G)$, we let $G[X,Y]$ denote the subgraph of $G$ with the vertex set $X\cup Y$, whose edges are the edges of $G$ with one end in $X$ and the other end in $Y$, and we let $e_{G}(X,Y)$ denote the number of edges in $G[X,Y]$. If $X=Y$, for convenience, we write $G[X]$ instead of $G[X,X]$, and $e_{G}(X)$ instead of $e_{G}(X,X)$. The maximum degree of a graph $G$ is denoted by $\Delta(G)$. For a vertex $v\in V(G)$, let $N_{G}(v)$ denote the set of neighbours of $v$ in $G$. For a set $S\subseteq V(G)$ and $v\in V(G)$, let $d_{G}(v,S)$ denote the number of edges in $G$ between $v$ and $S$. If the host graph is clear from the context, sometimes we drop the subscript and write $e(X,Y),e(X),d(v,S),N(v)$ instead of $e_{G}(X,Y),e_{G}(X),d_{G}(v,S),N_{G}(v)$ respectively. For a set $U\subseteq V(G)$, let $\overline{U}$ denote the set V(G)UV(G)-U, and let GUG-U denote the subgraph of $G$ induced by V(G)UV(G)-U.
For a graph $G$, a cut of $G$ is a partition $\{X,Y\}$ of $V(G)$, where $X$ and $Y$ are both non-empty. We say that a cut $\{X,Y\}$ is $\alpha$-sparse if $e_{G}(X,Y)\leq\alpha|X||Y|$. If $G$ is an $n$-vertex graph, we say that $G$ is $\alpha$-almost-bipartite if there exists a partition $\{X,Y\}$ of $V(G)$ such that $G$ has at most $\alpha n^{2}$ edges that are not in $G[X,Y]$. Otherwise we say that $G$ is $\alpha$-far-from-bipartite.
For a graph $G$ and vertices $x,y$ in $G$, an $(x,y)$-path in $G$ is a path with ends $x$ and $y$. If $P$ is an $(x,y)$-path and $Q$ is a $(y,z)$-path, we denote by $xPyQz$ the concatenation of the paths $P$ and $Q$.
Using a slight abuse of notation, for a collection $\mathcal{K}$ of subgraphs in a graph $G$, we denote the set of vertices covered by these subgraphs as $V(\mathcal{K})$. Throughout this paper, we often omit floor and ceiling signs when dealing with large numbers whenever they are not crucial.
2 Sketch of the proof of Theorem 1.4
In this section, we present a sketch of the proof of our main result, Theorem 1.4. At a high level, our approach begins by partitioning the vertex set of the regular graph $G$ into subsets with certain expansion properties, which we refer to as expanders. We then move some vertices between these expanders and remove a carefully constructed collection of $K_{t,t}$’s to ‘balance’ these expanders, using our first key lemma (Lemma 4.1). More precisely, Lemma 4.1 allows us to transform expanders that are bipartite (or close to bipartite) with imbalanced part sizes into balanced bipartite expanders. Next, we find almost perfect $K_{t,t}$-packings in each of the resulting expanders with the help of our second key lemma (Lemma 5.1). Several significant challenges arise in implementing this approach, which we now address in detail.
Let $F$ be a bipartite graph, and let $0<c\leq 1$. Let $G$ be a $d$-regular graph of order $n$, where $d\geq cn$. Let $t=|V(F)|$. Recall that Theorem 1.4 aims to find an $F$-packing that covers all but at most a constant number of vertices of $G$. It is easy to see that in order to find such a packing, it suffices to find a $K_{t,t}$-packing in $G$ covering all but at most $C$ vertices of $G$ for some constant $C=C(t,c)$, because $K_{t,t}$ has a perfect $F$-packing.
Our proof begins by partitioning the vertex set of the regular graph $G$ into a small number of subsets $Z_{1},\ldots,Z_{r}$ (for some $r\leq\lceil 1/c\rceil$) such that, for each $i\in[r]$, the induced subgraph $G[Z_{i}]$ has strong expansion properties, is either almost bipartite or far from bipartite, and the number of edges of $G$ between different sets $Z_{i}$ is small. (See Lemma 3.2 for the precise statement.)
2.1 Balancing the expanders
Given the partition $\{Z_{1},\ldots,Z_{r}\}$ of the vertices of $G$, a natural strategy for obtaining an almost perfect $K_{t,t}$-packing in $G$ is to find an almost perfect $K_{t,t}$-packing in each of the expanders $G[Z_{i}]$. However, it is possible that some of these expanders are rather imbalanced bipartite graphs, rendering this plan impossible. To overcome this obstacle, in our first key lemma (Lemma 4.1), we modify the sets $Z_{i}$ and remove a small collection of vertex-disjoint $K_{t,t}$-copies and a very small number of additional vertices so that the remainder of each modified set $Z_{i}$ spans an expander that is either bipartite and balanced or is far from bipartite. (In other words, this ‘balances’ the expanders $G[Z_{i}]$ that are almost bipartite.)
The first idea towards balancing the almost bipartite expanders is to use the regularity of $G$. For the rest of this section, let us assume for simplicity that $G$ itself is bipartite, and denote the bipartition of $G$ by $\{X,Y\}$ and the corresponding bipartition of $Z_{i}$ by $\{X_{i},Y_{i}\}$. Write $V:=V(G)$, and let $H$ denote the subgraph of $G$ on vertex set $V$ consisting of all edges whose endpoints lie in distinct sets $Z_{i}$.
A simple counting argument, using that $G$ is $d$-regular, shows that the following holds for every $i\in[r]$.
$e_{H}(X_{i},V-Z_{i})-e_{H}(Y_{i},V-Z_{i})=e_{G}(X_{i},V-Z_{i})-e_{G}(Y_{i},V-Z_{i})=(|X_{i}|-|Y_{i}|)\cdot d.$ (1)
Pretending for a moment that $t=1$, so that a $K_{t,t}$-packing is just a matching, a natural aim is thus to find a matching consisting of $\frac{1}{d}e_{H}(X_{i},V-Z_{i})$ edges leaving $Z_{i}$ and incident to $X_{i}$, and $\frac{1}{d}e_{H}(Y_{i},V-Z_{i})$ edges leaving $Z_{i}$ and incident to $Y_{i}$. By (1), this guarantees that the matching leaves the same number of vertices uncovered in $X_{i}$ and $Y_{i}$. [^1] We ignore rounding errors due to taking ceilings or floors; this is justified as we are allowed $O(1)$ uncovered vertices. To achieve this simultaneously for all $i\in[r]$, it suffices to find a matching consisting of $\frac{1}{d}e_{H}(X_{i},Y_{j})$ edges between $X_{i}$ and $Y_{j}$, for every distinct $i,j\in[r]$.
It turns out that it is quite easy to construct such a matching if all bipartite subgraphs $H[X_{i},Y_{j}]$ for $i\neq j$ have maximum degree at most $\rho n$, for some small constant $0<\rho\ll c$. Indeed, we consider the pairs $(X_{i},Y_{j})$ with $i\neq j$ one by one, in increasing order of $e_{H}(X_{i},Y_{j})$, and at each step we construct a matching of size $\frac{1}{d}e_{H}(X_{i},Y_{j})$ in $H[X_{i},Y_{j}]$ that is vertex-disjoint from all previously defined matchings. We claim that this is always possible. Indeed, consider a pair $(X_{i},Y_{j})$. Denote by $e$ the number of edges between $X_{i}$ and $Y_{j}$ and by $X_{i}^{\prime}\subseteq X_{i}$ and $Y_{j}^{\prime}\subseteq Y_{j}$ the sets of vertices that were used in previously defined matchings. Then $|X_{i}^{\prime}|,|Y_{j}^{\prime}|\leq r^{2}\cdot e/d$, by the order in which we process the pairs. By the maximum degree assumption and since $d\geq cn$, we have
$e_{H}(X_{i}-X_{i}^{\prime},Y_{j}-Y_{j}^{\prime})\geq e_{H}(X_{i},Y_{j})-\rho n\cdot(|X_{i}^{\prime}|+|Y_{j}^{\prime}|)\geq e-e/2=e/2.$
Since every graph $H^{\prime}$ has a matching of size at least $e(H^{\prime})/2\Delta(H^{\prime})$, it follows that there is a matching of size at least $e/(4\rho n)\geq e/d$ in $H[X_{i}-X_{i}^{\prime},,Y_{j}-Y_{j}^{\prime}]$, as required.
For general $t\geq 1$, our aim is the same, except that we replace matching edges between $X_{i}$ and $Y_{j}$ with copies of $K_{t,t}$ in $G$, each consisting of one vertex in $X_{i}$, $t$ vertices in $Y_{j}$, and t1t-1 vertices in $X_{j}$, so as to achieve the same balancing effect as moving one vertex from $X_{i}$ to $X_{j}$. The above greedy algorithm still applies if we use an asymmetric version of the Kővári–Sós–Turán theorem (Lemma 3.8) for all pairs $(X_{i},Y_{j})$ such that $H[X_{i},Y_{j}]$ contains at least $Cn$ edges, where $C$ is a large constant. Since we allow $O(1)$ uncovered vertices, we may ignore pairs $(X_{i},Y_{j})$ with fewer edges. Indeed, this yields a ‘balancing’ $K_{t,t}$-packing $\mathcal{K}$, which satisfies $|X_{i}-V(\mathcal{K})|=|Y_{i}-V(\mathcal{K})|+O(1)$ for all $i$.
The main challenge in proving Lemma 4.1 is therefore to emulate the situation in which the maximum degree (between the sets $Z_{i}$) is bounded by $\rho n$, which we achieve as follows. We perform the procedure below, sequentially modifying the graph $H$ and the sets $Z_{1},\ldots,Z_{r}$. If there is a pair $(X_{i},Y_{j})$ with $e_{H}(X_{i},Y_{j})\geq d$ and a vertex $v\in X_{i}$ that has at least $\rho n$ neighbours in $H$ lying in $Y_{j}$, then we perform the following three operations: we move $v$ from $X_{i}$ to $X_{j}$; we remove all edges between $v$ and $Y_{j}$ from $H$; and we remove $d-d_{H}(v)$ edges between $X_{i}-{v}$ and $Y_{j}$ from $H$, chosen arbitrarily. (The fact that the third operation is possible follows from the assumption that $e_{H}(X_{i},Y_{j})\geq d$.) Note that this procedure only removes edges from $H$ in each step and never adds any. In particular, although $v$ is now moved to $X_{j}$ we do not add to $H$ the edges from $v$ to $Y_{i}$. This shows that the number of edges in $H$ decreases by at least $\rho n$ in each step of the procedure. The first two operations are quite natural: since $v$ has large degree into $Y_{j}$, adding it to $X_{j}$ would preserve the expansion properties of $G[X_{j},Y_{j}]$. Since $H$ initially contained no edges between $X_{\ell}$ and $Y_{\ell}$ for any $\ell$, the second operation preserves this condition after each step of the procedure. The motivation for the third operation is the observation that if $H,X_{1},\ldots,X_{r},Y_{1},\ldots,Y_{r}$ satisfy (1) and we modify them by performing the three operations above, then the resulting $H,X_{1},\ldots,X_{r},Y_{1},\ldots,Y_{r}$ still satisfy (1). Indeed, for example, to see that (1) continues to hold for $X_{i},Y_{i}$, note that after performing the three operations above, both the left-hand and right-hand sides of (1) decrease by exactly $d$.
Continuing this procedure for as long as possible (also with the roles of $X_{i}$ and $Y_{i}$ reversed), the resulting graph $H$ and sets $X_{1},\ldots,X_{r},Y_{1},\ldots,Y_{r}$ still satisfy (1), and for every $i\neq j$, the bipartite subgraph $H[X_{i},Y_{j}]$ either has maximum degree at most $\rho n$ or contains fewer than $d$ edges. Therefore, this allows us to run the above procedure on $H$ to find a $K_{t,t}$-packing $\mathcal{K}$ that balances the modified sets $Z_{i}$ so that $|X_{i}-V(\mathcal{K})|=|Y_{i}-V(\mathcal{K})|+O(1)$ for every $i$. Moreover, since $H$ initially has few edges (as there are few edges between the sets $Z_{i}$) and each step removes at least $\rho n$ edges, we are guaranteed that only few vertices are moved throughout the procedure; this is essential for showing that the modified sets $Z_{i}$ are still expanders.
Balancing expanders (in different contexts) was also a key step in the proofs of Kühn, Lo, Osthus and Staden in [27, 28] and of Gruslys and Letzter in [11]. The method we develop here to construct a balanced $K_{t,t}$-packing differs significantly from these previous approaches.
2.2 Packing $K_{t,t}$’s in expanders
Having balanced the expanders (using Lemma 4.1), the original problem is reduced to the problem of finding an almost perfect $K_{t,t}$-packing (covering all but a constant number of vertices) in an expander $G$ that is dense, almost regular, and either far from bipartite or balanced and bipartite. We address this reduced problem in our second key lemma, Lemma 5.1, where we construct a perfect $K_{t,t}$-packing in such an expander under the assumption that the number of vertices is divisible by $2t$.
The proof of Lemma 5.1 uses the regularity lemma and combines two main themes. First, we exploit the fact that dense, almost regular expanders that are either far from bipartite or bipartite and balanced contain Hamilton cycles, following the work of Kühn, Osthus, and Treglown [34] on robust expanders. Second, as in the previous section, we again use $K_{t,t}$-packings for balancing purposes, but now in two distinct ways that differ from the approach used earlier.
To begin, we aim to find disjoint sets $U_{1},\ldots,U_{2m}$ of equal size that cover almost all vertices of $G$, such that almost every pair $(U_{i},U_{j})$ is $\varepsilon$-regular in $G$, and each pair $G[U_{2i-1},U_{2i}]$ is somewhat dense and super-regular (that is, almost every pair $(U_{i},U_{j})$ is $\varepsilon$-regular in $G$, and moreover $G[U_{2i-1},U_{2i}]$ has reasonably large minimum degree). To achieve this, we apply the regularity lemma (Lemma 3.5) to $G$ to obtain vertex-disjoint clusters $V_{1},\ldots,V_{m}$ of equal size that cover almost all vertices of $G$ and such that almost all pairs are $\varepsilon$-regular. We then use the expansion properties of $G$, together with the aforementioned result on Hamiltonicity of expanders, to deduce that the reduced graph $\Gamma$ (whose vertices are $V_{1},\ldots,V_{m}$ and whose edges correspond to dense $\varepsilon$-regular pairs) admits a perfect fractional matching. This, in turn, implies the existence of a perfect $2$-matching in $\Gamma$, that is, a collection of vertex-disjoint edges and cycles covering all vertices. Consequently, splitting each cluster $V_{i}$ into two equal parts yields a perfect matching in the corresponding reduced graph. By possibly removing a small proportion of vertices from each of these new clusters, we can ensure that each matched pair is super-regular, as required. The idea of seeking perfect $2$-matchings in reduced graphs has appeared previously, for example in [17, 16, 25].
Having found the sets $U_{1},\ldots,U_{2m}$, we can apply the blow-up lemma to each pair $G[U_{2i-1},U_{2i}]$ to cover all but $2t$ vertices of $U_{2i-1}\cup U_{2i}$ with a $K_{t,t}$-packing. However, this approach may only produce a $K_{t,t}$-packing in $G$ that leaves linearly many vertices uncovered, namely those outside the sets $U_{1},\ldots,U_{2m}$. To cover all vertices of $G$ (rather than almost all), we instead find a $K_{t,t}$-packing $\mathcal{K}$ that covers every vertex outside $U_{1}\cup\cdots\cup U_{2m}$ and only very few vertices in each $U_{i}$, and, crucially, leaves the same number of uncovered vertices in each $U_{i}$, where this number is divisible by $t$. Once such a $K_{t,t}$-packing $\mathcal{K}$ is found, we can apply the blow-up lemma to cover the remaining vertices in each pair $G[U_{2i-1},U_{2i}]$ with a perfect $K_{t,t}$-packing, thereby completing $\mathcal{K}$ to a perfect $K_{t,t}$-packing in $G$.
To construct such a $K_{t,t}$-packing $\mathcal{K}$, we proceed in three steps. First, we construct a packing $\mathcal{K}_{1}$ that covers all vertices outside $U_{1}\cup\cdots\cup U_{2m}$ while using very few vertices from each set $U_{i}$. This is straightforward using the Kővári–Sós–Turán theorem together with the properties of $G$. Next, we construct two further $K_{t,t}$-packings, $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$, each serving a different balancing role. The packing $\mathcal{K}_{2}$ is chosen so that, for every $i$, the number of vertices in $U_{i}$ uncovered by $\mathcal{K}_{1}\cup\mathcal{K}_{2}$ is divisible by $t$. We then construct $\mathcal{K}_{3}$ so that the number of vertices in each $U_{i}$ uncovered by $\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}$ is the same for all $i$, and still divisible by $t$. Taking $\mathcal{K}:=\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3}$ allows us to complete the proof using the blow-up lemma, as explained in the previous paragraph.
For the construction of $\mathcal{K}_{2}$, the key idea is that a walk $U_{i_{1}}\ldots U_{i_{2s-1}}$ of even length in the reduced graph allows us to “transfer” vertices from $U_{i_{1}}$ to $U_{i_{2s-1}}$ (modulo $t$). More precisely, we construct a collection of copies of $K_{t,t}$ between the clusters corresponding to this walk so as to create a balancing effect on the clusters at the two ends of the walk when these copies are removed. By combining several such walks, we can construct a collection $\mathcal{K}_{2}$ of copies of $K_{t,t}$ such that the number of uncovered vertices in each $U_{i}$ is divisible by $t$, as desired (see Section 5.4 for further details).
We now describe the construction of $\mathcal{K}_{3}$. For each $i\in[2m]$, let $U_{i}^{\prime}$ denote the set of vertices in $U_{i}$ that remain uncovered. Our first step is to find a template matching $\mathcal{M}$ in $G[U_{1}^{\prime}\cup\ldots\cup U_{2m}^{\prime}]$ such that $|U_{i}^{\prime}|-t\cdot|V(\mathcal{M})\cap U_{i}^{\prime}|$ is the same for all $i\in[2m]$. Given such a matching $\mathcal{M}$, we apply the blow-up lemma to construct a $K_{t,t}$-packing $\mathcal{K}_{3}$ by replacing each edge of $\mathcal{M}$ between $U_{i}^{\prime}$ and $U_{j}^{\prime}$ with a copy of $K_{t,t}$ in $G[U_{i}^{\prime},U_{j}^{\prime}]$. Consequently, the number of vertices in $U_{i}^{\prime}$ that remain uncovered by $\mathcal{K}_{3}$ is exactly $|U_{i}^{\prime}|-t\cdot|V(\mathcal{M})\cap U_{i}^{\prime}|$. By the choice of $\mathcal{M}$, this quantity is the same for all $i\in[2m]$; moreover, it is divisible by $t$ by choice of $\mathcal{K}_{2}$, as required. Finally, to construct the template matching $\mathcal{M}$, we select subsets $U_{i}^{\prime\prime}\subseteq U_{i}$ of appropriate size and show that the induced graph $G[U_{1}^{\prime\prime}\cup\ldots\cup U_{2m}^{\prime\prime}]$ inherits the relevant expansion properties of $G$. By the aforementioned Hamiltonicity of expanders, this graph contains a perfect matching, which serves as the desired template matching $\mathcal{M}$ (see Section 5.5 for further details).
3 Useful lemmas and tools
3.1 Concentration inequality
We need the following concentration inequality for the hypergeometric distribution. Let $m,n,N\in\mathbb{N}$ such that $\max\{m,n\}<N$. Recall that a random variable $X$ has hypergeometric distribution with parameters $N,n,m$ if $X\coloneqq|S\cap[m]|$, where $S$ is a random subset of $[N]$ of size $n$.
Lemma 3.1 (Theorem 2.10 in [14]).
Let $X\sim\operatorname{Hypergeo}(N,m,n)$ be a hypergeometric random variable with mean $\mathbb{E}[X]=nm/N$. Then for all $0\leq\varepsilon\leq 3/2$, we have
$\mathbb{P}\big(|X-\mathbb{E}[X]|\geq\varepsilon\mathbb{E}[X]\big)\leq 2e^{-\varepsilon^{2}\mathbb{E}[X]/3}.$
3.2 Expanders in dense regular graphs
We will use the following structural result (Lemma 3.2) due to Gruslys and Letzter [11], which goes back to the work of Kühn, Lo, Osthus, and Staden [27, 28]. This result serves as the first step in the proofs of both Theorem 1.4 and Theorem 1.7. It shows that the vertices of a dense regular graph $G$ can be partitioned into a small number of sets such that the subgraphs induced by these sets exhibit strong expansion properties and have a large minimum degree. Throughout the remainder of the paper, we refer to these subgraphs as expanders. The notion of expansion used here is characterized by the absence of sparse cuts. Specifically, a graph $G$ has no $\zeta$-sparse cuts if, for every partition $\{X,Y\}$ of $V(G)$, the number of edges between $X$ and $Y$ is at least $\zeta|X||Y|$. This notion of expansion was first introduced (using different terminology and with $\zeta$ as a function of $n$) in a paper by Conlon, Fox, and Sudakov [7]. We will distinguish between expanders that are close to bipartite and those that are far from bipartite. Recall that a graph $G$ on $n$ vertices is said to be $\gamma$-almost-bipartite if one can make $G$ bipartite by removing up to $\gamma n^{2}$ edges, and otherwise, $G$ is called $\gamma$-far-from-bipartite.
Lemma 3.2 (Lemma 2.1 from [11]).
Let $c\in(0,1)$ and $n_{0}\in\mathbb{N}$ satisfy $1/n_{0}\ll c$. Let $G$ be a $d$-regular graph on $n$ vertices, where $n\geq n_{0}$ and $d\geq cn$. Then, there exist positive numbers $r\leq\left\lceil 1/c\right\rceil$ and $\eta,\beta,\gamma,\zeta,\delta$, where $1/n_{0}\ll\eta\ll\beta\ll\gamma\ll\zeta\ll\delta\ll c$, and a partition $\{Z_{1},\ldots,Z_{r}\}$ of $V(G)$ that satisfies the following properties.
  1. (E1)
    $G$ has at most $\eta n^{2}$ edges with ends in distinct $Z_{i}$’s.
  2. (E2)
    $G[Z_{i}]$ has minimum degree at least $\delta n$, for $i\in[r]$.
  3. (E3)
    $G[Z_{i}]$ has no $\zeta$-sparse cuts, for $i\in[r]$.
  4. (E4)
    $G[Z_{i}]$ is either $\beta$-almost-bipartite or $\gamma$-far-from-bipartite, for $i\in[r]$.
The following lemma is used multiple times throughout the paper. Its proof is very similar to that of Lemma 2.2 in [11], and relies on results concerning Hamiltonicity of robust expanders, due to Kühn, Osthus and Treglown [34]. The lemma shows that expanders far from being bipartite are Hamiltonian and remain Hamiltonian even after the removal of any small set of vertices. However, expanders that are bipartite may not be Hamiltonian – for example, an imbalanced bipartite graph cannot be Hamiltonian. In this case, the lemma shows that such expanders still become Hamiltonian after the removal of any small set of vertices that balances its two parts. We prove Lemma 3.3 in Appendix B.
Lemma 3.3.
Let $1/n\ll\eta\ll\xi\ll\gamma,\zeta\ll c$, and let $d\geq cn$. Suppose $G$ is an $n$-vertex graph with maximum degree at most $d$ and average degree at least $d-\eta n$, such that $G$ does not have $\zeta$-sparse cuts. Additionally, assume that $G$ is bipartite or $\gamma$-far-from-bipartite. Then the following holds.
  1. (HC1)
    If $G$ is bipartite with the bipartition $\{X,Y\}$, then, for every subset $W\subseteq V(G)$ of size at most $\xi n$ satisfying |XW|=|YW||X-W|=|Y-W| and any vertices $x\in X-W$, $y\in Y-W$, there is a Hamilton path in GWG-W with ends $x$ and $y$. In particular, GWG-W has a Hamilton cycle.
  2. (HC2)
    If $G$ is $\gamma$-far-from-bipartite, then, for every subset $W\subseteq V(G)$ of size at most $\xi n$ and any two distinct vertices $x,y\in V(G)-W$, there is a Hamilton path in GWG-W with ends $x$ and $y$. In particular, GWG-W has a Hamilton cycle.
We use the following lemma from [11] in the proof of Theorem 1.7. When presented with a partition of a dense regular graph into expanders (by applying Lemma 3.2), Lemma 3.4 produces a small linear forest (i.e., a small collection of vertex-disjoint paths) whose removal balances the expanders that are close to being bipartite. A similar linear forest was also constructed in [27, 28].
Lemma 3.4 (Lemma 2.3 from [11]).
Let $\eta,\beta,\xi,\gamma,\zeta,\delta,c\in(0,1)$ and let $n\in\mathbb{N}$ satisfy $1/n\ll\eta\ll\beta\ll\xi\ll\gamma\ll\zeta\ll\delta\ll c$. Let $G$ be a $d$-regular graph on $n$ vertices, where $d\geq cn$. Suppose that $\{Z_{1},\ldots,Z_{r}\}$ is a partition of $V(G)$ satisfying properties (E1)(E4) in Lemma 3.2, where $r\leq\left\lceil 1/c\right\rceil$. For $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite, let $\{X_{i},Y_{i}\}$ be a partition of $Z_{i}$ maximising $e_{G}(X_{i},Y_{i})$. Then there is a linear forest $H$ in $G$ with the following properties.
  1. (P1)
    $|V(H)|\leq\xi n$.
  2. (P2)
    $H$ has no isolated vertices.
  3. (P3)
    For each $i\in[r]$, $Z_{i}$ contains either zero or two leaves of $H$. Moreover, for each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite and $Z_{i}$ contains two leaves of $H$, one of the leaves is in $X_{i}$ and the other leaf is in $Y_{i}$.
  4. (P4)
    For each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite, $|X_{i}-V(H)|=|Y_{i}-V(H)|$.
3.3 The regularity lemma and the blow-up lemma
Given a graph $G$, and sets $X,Y\subseteq V(G)$, let $d(X,Y)\coloneqq\frac{e(X,Y)}{|X||Y|}$. The density of a bipartite graph $G$ with the bipartition $\{A,B\}$ is denoted by $d(A,B).$
Given $\varepsilon>0$, we say that a bipartite graph $G$ with the bipartition $\{A,B\}$ is $\varepsilon$-regular if, for all sets $X\subseteq A$ and $Y\subseteq B$ with $|X|\geq\varepsilon|A|$ and $|Y|\geq\varepsilon|B|$, we have $|d(A,B)-d(X,Y)|<\varepsilon.$ The following is the degree form of Szemerédi’s Regularity Lemma which can be easily derived from the classical version (see, e.g., [3] and [8]). It is a slight variation of Lemma 2.4 in [31].
Lemma 3.5 (regularity lemma).
For every $\varepsilon>0$, there exists $M=M(\varepsilon)$ such that the following holds. Let $0\leq\mu\leq 1$, and let $G$ be an $n$-vertex graph. Then there is a partition $\{V_{0},\ldots,V_{m}\}$ of $V(G)$ and a spanning subgraph $G^{\prime}$ of $G-V_{0}$, such that the following properties hold.
  1. (R1)
    $m\leq M$.
  2. (R2)
    If $G$ is bipartite, then every $V_{i}$ with $i\in[m]$ is contained in one of the two parts of $G$.
  3. (R3)
    $|V_{0}|\leq\varepsilon n$ and $|V_{1}|=\ldots=|V_{m}|\leq\varepsilon n$.
  4. (R4)
    $d_{G^{\prime}}(v)\geq d_{G}(v)-(\mu+\varepsilon)n$ for every $v\in V(G^{\prime})$.
  5. (R5)
    $e_{G^{\prime}}(V_{i})=0$ for every $i\in[m]$.
  6. (R6)
    For all $1\leq i,j\leq m$, the graph $G^{\prime}[V_{i},V_{j}]$ is $\varepsilon$-regular with density either 0 or more than $\mu$.
The sets $V_{1},\ldots,V_{m}$ in Lemma 3.5 are called clusters, and the set $V_{0}$ is called the exceptional set. Given clusters and $G^{\prime}$ as in Lemma 3.5, the reduced graph $\Gamma$ is the graph whose vertices are $V_{1},\ldots,V_{m}$ (note that $V_{0}$ is omitted here) and in which $V_{i}$ is adjacent to $V_{j}$ whenever $G^{\prime}[V_{i},V_{j}]$ is $\varepsilon$-regular and has density more than $\mu$. Thus, $V_{i}V_{j}$ is an edge of $\Gamma$ if and only if $G^{\prime}$ has an edge between $V_{i}$ and $V_{j}$.
Given $\varepsilon>0$ and $d\in[0,1]$, a bipartite graph $G$ with the bipartition $\{A,B\}$ is called $(\varepsilon,d)$-super-regular, if all sets $X\subseteq A$ and $Y\subseteq B$ with $|X|\geq\varepsilon|A|$ and $|Y|\geq\varepsilon|B|$ satisfy $d(X,Y)>d$, and moreover, every vertex in $A$ has degree more than $d|B|$, and every vertex in $B$ has degree more than $d|A|$.
We will need the following special case of the blow-up lemma of Komlós, Sárközy and Szemerédi [21].
Lemma 3.6 (Special case of the blow-up lemma [21]).
For every $d>0$ and $\Delta$, there exists a positive number $\varepsilon_{0}=\varepsilon_{0}(d,\Delta)$ such that for all $\varepsilon\leq\varepsilon_{0}$ the following holds. If $H$ is a balanced complete bipartite graph with $n$ vertices in each part satisfying $\Delta(H)\leq\Delta$, then every $(\varepsilon,d)$-super-regular bipartite graph with $n$ vertices in each part contains $H$ as a subgraph.
We also need the following standard fact concerning super-regular graphs.
Proposition 3.7 (Proposition 2.3 in [31]).
Every $\varepsilon$-regular bipartite graph $G$ with the bipartition $\{A,B\}$ and density $d>2\varepsilon$, can be made into a $(\varepsilon/(1-\varepsilon),d-2\varepsilon)$-super-regular graph by removing $\varepsilon|A|$ vertices from $A$ and $\varepsilon|B|$ vertices from $B$.
3.4 The Kővári–Sós–Turán theorem
The following is an unbalanced variant of the well-known Kővári–Sós–Turán theorem [26].
Lemma 3.8.
Let $t\geq 2$, let $\delta>0$, let $G$ be a bipartite graph on at most $n$ vertices with the bipartition $\{X,Y\}$ such that $|X|\geq 2t\left(\frac{e}{\delta}\right)^{t}$ and every vertex in $X$ has at least $\delta n$ neighbours in $Y$. Then $G$ contains a $K_{t,t}$.
Suppose for a contradiction that $G$ is $K_{t,t}$-free. A copy of $K_{1,t}$ in $G$ is called a $t$-star, the vertex of degree $t$ in a $t$-star is called its center, and the $t$ vertices of degree $1$ in a $t$-star are called its leaves. Since $G$ is $K_{t,t}$-free, the number of $t$-stars in $G$ whose centre is in $X$ and whose $t$ leaves are in $Y$ is at most
$(t-1)\binom{|Y|}{t}\leq t\left(\frac{e|Y|}{t}\right)^{t}\leq\frac{(en)^{t}}{t^{t-1}}.$
On the other hand, since every vertex in $X$ has at least $\delta n$ neighbours in $Y$, the number of such $t$-stars in $G$ is at least
$|X|\binom{\delta n}{t}\geq|X|\left(\frac{\delta n}{t}\right)^{t}.$
Thus, we obtain $|X|\leq t\left(\frac{e}{\delta}\right)^{t}.$ This is a contradiction, proving the lemma. ∎
4 Balancing the expanders
As discussed in the proof sketch, our proof of Theorem 1.4 starts by partitioning the vertex set of our dense regular graph $G$ into a small number of expanders (using Lemma 3.2). A natural strategy for finding a $K_{t,t}$-packing in $G$ is to find a $K_{t,t}$-packing in each of these expanders. However, if any of these expanders happens to be a rather imbalanced bipartite graph, then it is impossible to find a $K_{t,t}$-packing in it covering all but a constant number of vertices. To overcome this difficulty, we need the following key lemma which shows that we can partition the vertex set of any dense regular graph $G$ into a set $L$ that contains a perfect $K_{t,t}$-packing, a small set of constant size, and a small number of expanders, each of which is either far from bipartite or bipartite and balanced. Using Corollary 5.2 we can then find a $K_{t,t}$-packing in each of these expanders covering all but a constant number of its vertices, giving us the desired $K_{t,t}$-packing in $G$.
Lemma 4.1.
Let $0<c\leq 1$, let $t\geq 2$ be an integer, and let $n_{0}$ be sufficiently large. Suppose that $G$ is a $d$-regular graph on $n$ vertices, where $d\geq cn$ and $n\geq n_{0}$. Then there exist positive numbers $\eta,\gamma,\zeta,r$, with $r\leq\left\lceil 1/c\right\rceil$ and $1/n_{0}\ll\eta\ll\zeta\ll\gamma\ll c,1/t$ such that the following holds. There exist pairwise disjoint sets $Z_{1},\ldots,Z_{r},L\subseteq V(G)$ and a subgraph $G^{\prime}\subseteq G$ with $V(G^{\prime})=V(G)$ satisfying the following properties.
  1. (B1)
    $|V(G)-(Z_{1}\cup\ldots\cup Z_{r}\cup L)|\leq 64r^{2}\cdot\frac{t}{c}\cdot\left(\frac{e}{8\zeta}\right)^{t}$.
  2. (B2)
    For every $i\in[r]$, $G^{\prime}[Z_{i}]$ has average degree at least $d-\eta|Z_{i}|$.
  3. (B3)
    For every $i\in[r]$, $G^{\prime}[Z_{i}]$ has no $\zeta$-sparse cuts.
  4. (B4)
    For every $i\in[r]$, $G^{\prime}[Z_{i}]$ is either bipartite and balanced, or one needs to remove at least $\gamma n^{2}$ edges from $G^{\prime}[Z_{i}]$ to make it bipartite.
  5. (B5)
    $G[L]$ has a perfect $K_{t,t}$-packing.
Let $c,t,n_{0},n,d$ be as in the statement of the lemma, and let $G$ be a $d$-regular graph on $n$ vertices, where $d\geq cn$. By applying Lemma 3.2 to $G$ (with $\min\{c,1/t\}$ playing the role of $c$) we obtain positive numbers $r,\eta,\beta,\gamma,\zeta,\delta$, and a partition $\{Z_{1},\ldots,Z_{r}\}$ of $V(G)$ satisfying $r\leq\left\lceil 1/c\right\rceil$ and (E1)(E4), such that
$1/n_{0}\ll\eta\ll\beta\ll\gamma\ll\zeta\ll\delta\ll c,1/t.$
Let $\sigma,\rho$ be positive numbers satisfying
$\beta\ll\sigma\ll\rho\ll\gamma.$
For every $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite, fix $\{X_{i},Y_{i}\}$ to be a partition of $Z_{i}$ that maximises the number of edges between $X_{i}$ and $Y_{i}$. For every $i\in[r]$ such that $G[Z_{i}]$ is $\gamma$-far-from-bipartite, let $\{X_{i},Y_{i}\}$ be a partition of $Z_{i}$ satisfying
  1. (G1)
    $|e_{G}(X_{i})-e_{G}(Y_{i})|\leq\beta n^{2}$,
  2. (G2)
    $\delta(G[X_{i},Y_{i}])\geq\delta n/3$.
Notice that this can easily be achieved by taking $\{X_{i},Y_{i}\}$ to be a random partition of $Z_{i}$. Moreover, these conditions are satisfied also when $G[Z_{i}]$ is $\beta$-almost-bipartite. Indeed, if $G[Z_{i}]$ is $\beta$-almost-bipartite, then $e_{G}(X_{i})+e_{G}(Y_{i})\leq\beta n^{2}$, implying that (G1) holds. Moreover, since $\{X_{i},Y_{i}\}$ is a partition maximising the number of edges in $G[X_{i},Y_{i}]$ and since the minimum degree in $G[Z_{i}]$ is at least $\delta n$ (by (E2)), the graph $G[X_{i},Y_{i}]$ has minimum degree at least $\delta n/2$, showing (G2).
As mentioned in the proof sketch, we would like to use the regularity of $G$ to correct the imbalance between $X_{i}$ and $Y_{i}$, for every $i\in[r]$. This will be done via the following equation, which we claim holds for every $i\in[r]$, and which generalises (1) from the proof sketch from bipartite graphs to general graphs.
$\left(e_{G}(X_{i},\overline{Z_{i}})-e_{G}(Y_{i},\overline{Z_{i}})\right)+2\left(e_{G}(X_{i})-e_{G}(Y_{i})\right)-d\left(|X_{i}|-|Y_{i}|\right)=0.$ (2)
Indeed, to see that (2) holds, we have the following equations counting the number of edges incident to $X_{i}$ and $Y_{i}$, respectively.
$d|X_{i}|$ $=e_{G}(X_{i},Y_{i})+e_{G}(X_{i},\overline{Z_{i}})+2e_{G}(X_{i}),$
$d|Y_{i}|$ $=e_{G}(X_{i},Y_{i})+e_{G}(Y_{i},\overline{Z_{i}})+2e_{G}(Y_{i}).$
Subtracting one from the other yields (2).
As explained in the proof sketch (see also Claim 4.3 below), (2) shows that to find a balancing $K_{t,t}$-packing, it suffices to find disjoint $K_{t,t}$-packings for each pair $(W_{i},W_{j})$ with $W_{i}\in\{X_{i},Y_{i}\}$ and $W_{j}\in\{X_{j},Y_{j}\}$ of size roughly $\frac{1}{d}e_{G}(W_{i},W_{j})$, whose $K_{t,t}$’s effectively move one vertex from $W_{i}$ to $Z_{j}-W_{j}$. This can be done greedily if the maximum degree of edges between expanders is small (at most $\rho n$ for a small constant $\rho$). However, the maximum degree of edges between expanders need not be small, even if we are allowed to modify the $Z_{i}$’s somewhat. Nevertheless, we emulate the small maximum degree situation by constructing a subgraph $H$ of $G$, and modifying the sets $Z_{i}$, so that (2) is maintained and the maximum degree of the final $H$ is small (with possibly some, very few, exceptional vertices with large degree). To do so we first define $H$ as follows, and then modify it and the $Z_{i}$’s as described in the next subsection.
Let $H$ be the graph obtained from $G$ by removing all edges between $X_{i}$ and $Y_{i}$, and also arbitrarily removing $\min\{e_{G}(X_{i}),e_{G}(Y_{i})\}$ edges from each of $G[X_{i}]$ and $G[Y_{i}]$ for all $i\in[r]$. Then $H$ satisfies the following properties.
  1. (H1)
    $\left(e_{H}(X_{i},\overline{Z_{i}})-e_{H}(Y_{i},\overline{Z_{i}})\right)+2\left(e_{H}(X_{i})-e_{H}(Y_{i})\right)-d\left(|X_{i}|-|Y_{i}|\right)=0$,
  2. (H2)
    $e_{H}(X_{i},Y_{i})=0$ for every $i\in[r]$,
  3. (H3)
    $H$ has maximum degree at most $d-\delta n/3$,
  4. (H4)
    $e(H)\leq\eta n^{2}+r\beta n^{2}\leq 2r\beta n^{2}$.
Note that for (H3) we used (G2), and for (H4) we used (E1) and (G1).
In the next subsection, we use a procedure for moving vertices between the expanders $G[Z_{i}]$, $i\in[r]$, in order to ensure that the degrees of vertices in $H$ between the resulting subgraphs of $G$ are appropriately bounded (whenever there are sufficiently many such edges), while also ensuring that the subgraphs have large minimum degree. More precisely, we will show that, when the procedure ends, (H1)(H4) still hold, and (H5) and (H6) below also hold.
4.1 Procedure for moving vertices between the expanders for better degree control
Start with $H,(X_{i})_{i\in[r]},(Y_{i})_{i\in[r]},(Z_{i})_{i\in[r]}$ satisfying (H1)(H4), and modify them using the following procedure, by running it as long as possible. For convenience, in the rest of the proof we denote the values of these quantities before the start of the procedure by $H^{\textrm{start}}$, $(X^{\textrm{start}}_{i})_{i\in[r]}$, $(Y^{\textrm{start}}_{i})_{i\in[r]}$, $(Z^{\textrm{start}}_{i})_{i\in[r]}$, respectively.
1:procedure Move-high-degree-vertices
2:  Start with $H,(X_{i})_{i\in[r]},(Y_{i})_{i\in[r]},(Z_{i})_{i\in[r]}$ satisfying (H1)(H4).
3:  while there exist $i,j\in[r],W_{i}\in\{X_{i},Y_{i}\},W_{j}\in\{X_{j},Y_{j}\}$ and $v\in W_{i}$ such that
$d_{H}(v,W_{j})\geq\rho n\quad\text{ and }\quad e_{H}(W_{i},W_{j})\geq d.$ (3)
do
4:   $W_{i}\leftarrow W_{i}-\{v\}$,
5:   $U_{j}\leftarrow U_{j}\cup\{v\}$, where $\{U_{j},W_{j}\}=\{X_{j},Y_{j}\}$,
6:   $Z_{i}\leftarrow X_{i}\cup Y_{i}$ and $Z_{j}\leftarrow X_{j}\cup Y_{j}$,
7:   $H\leftarrow H-(E_{i,j}\cup E_{H}(v,W_{j}))$, where $E_{i,j}\subseteq H[W_{i}-\{v\},W_{j}-\{v\}]$ is a set of exactly $d-d_{H}(v)$ edges, and $E_{H}(v,W_{j})$ is the set of all edges in $H$ between $v$ and $W_{j}$.
8:  end while
9:end procedure
In other words, whenever there are $i,j\in[r]$, $W_{i}\in\{X_{i},Y_{i}\}$, $W_{j}\in\{X_{j},Y_{j}\}$ and $v\in W_{i}$ such that $d_{H}(v,W_{j})\geq\rho n$ and $e_{H}(W_{i},W_{j})\geq d$ (see (3)), we move $v$ from $W_{i}$ to $U_{j}$ (where $\{U_{j},W_{j}\}=\{X_{j},Y_{j}\}$), we remove $d-d_{H}(v)$ edges between $W_{i}-\{v\}$ and $W_{j}-\{v\}$ from $H$, and we also remove all edges between $v$ and $W_{j}$ from $H$. Note that it is possible to remove $d-d_{H}(v)$ edges between $W_{i}-\{v\}$ and $W_{j}-\{v\}$ by the assumption that $e_{H}(W_{i},W_{j})\geq d\geq d-d_{H}(v)+d_{H}(v,W_{j})$.
Claim 4.2.
Throughout the procedure Move-high-degree-vertices, (H1) to (H4) hold.
Since we never add any edges to $H$ during the procedure and (H3) and (H4) hold at the start, it is obvious that they hold throughout the procedure. It is also easy to check that (H2) is true, because it is true at the start, and after moving a vertex $v$ to $U_{j}$, we remove all edges between $v$ and $W_{j}$. Hence, it suffices to show that if (H1) holds for a given $H$, $(X_{i})_{i\in[r]}$, $(Y_{i})_{i\in[r]}$, then (H1) holds after performing one step of the procedure as well. Let $H^{\prime}$, $(X^{\prime}_{i})_{i\in[r]}$, $(Y^{\prime}_{i})_{i\in[r]}$ denote the corresponding quantities after performing one step of the procedure. We need to show that for all $i\in[r]$,
$(e_{H^{\prime}}(X^{\prime}_{i},\overline{Z^{\prime}_{i}})-e_{H^{\prime}}(Y^{\prime}_{i},\overline{Z^{\prime}_{i}}))+2\left(e_{H^{\prime}}(X^{\prime}_{i})-e_{H^{\prime}}(Y^{\prime}_{i})\right)-d\left(|X^{\prime}_{i}|-|Y^{\prime}_{i}|\right)=0.$ (4)
Without loss of generality, we first prove it when $(W_{i},W_{j})=(X_{i},Y_{j})$ with $i\neq j$ and then we prove it when $(W_{i},W_{j})=(X_{i},X_{i})$. Indeed, by symmetry, this covers all cases since $e_{H}(X_{i},Y_{i})=0$, and $X_{i}$ and $Y_{i}$ play the same role for every $i\in[r]$.
First, consider the case when $(W_{i},W_{j})=(X_{i},Y_{j})$ with $i\neq j$. In this case, note that the following twelve equations hold.
$e_{H^{\prime}}(X_{i}^{\prime},\overline{Z_{i}^{\prime}})-e_{H}(X_{i},\overline{Z_{i}})$ $=-(d-d_{H}(v))-d_{H}(v,\overline{Z_{i}})+d_{H}(v,X_{i})$
$e_{H^{\prime}}(Y_{i}^{\prime},\overline{Z_{i}^{\prime}})-e_{H}(Y_{i},\overline{Z_{i}})$ $=0$
$e_{H^{\prime}}(X_{i}^{\prime})-e_{H}(X_{i})$ $=-d_{H}(v,X_{i})$
$e_{H^{\prime}}(Y_{i}^{\prime})-e_{H}(Y_{i})$ $=0$
$|X_{i}^{\prime}|-|X_{i}|$ $=-1$
$|Y_{i}^{\prime}|-|Y_{i}|$ $=0.$
$e_{H^{\prime}}(X_{j}^{\prime},\overline{Z_{j}^{\prime}})-e_{H}(X_{j},\overline{Z_{j}})$ $=d_{H}(v,\overline{Z_{j}})-d_{H}(v,X_{j})$
$e_{H^{\prime}}(Y_{j}^{\prime},\overline{Z_{j}^{\prime}})-e_{H}(Y_{j},\overline{Z_{j}})$ $=-(d-d_{H}(v))-d_{H}(v,Y_{j})$
$e_{H^{\prime}}(X_{j}^{\prime})-e_{H}(X_{j})$ $=d_{H}(v,X_{j})$
$e_{H^{\prime}}(Y_{j}^{\prime})-e_{H}(Y_{j})$ $=0$
$|X_{j}^{\prime}|-|X_{j}|$ $=1$
$|Y_{j}^{\prime}|-|Y_{j}|$ $=0.$
Multiplying the first six equations by 1,1,2,2,d,d1,-1,2,-2,-$d,$ respectively, and adding them all up, we find that
$(e_{H^{\prime}}(X_{i}^{\prime},\overline{Z_{i}^{\prime}})-e_{H^{\prime}}(Y_{i}^{\prime},\overline{Z_{i}^{\prime}}))+2\left(e_{H^{\prime}}(X_{i}^{\prime})-e_{H^{\prime}}(Y_{i}^{\prime})\right)-d\left(|X_{i}^{\prime}|-|Y_{i}^{\prime}|\right)$
$=$ $\left(e_{H}(X_{i},\overline{Z_{i}}\right)-e_{H}(Y_{i},\overline{Z_{i}}))+2\left(e_{H}(X_{i})-e_{H}(Y_{i})\right)-d\left(|X_{i}|-|Y_{i}|\right)=0.$
Similarly, multiplying the last six equations by 1,1,2,2,d,d1,-1,2,-2,-$d,$ respectively, and adding them up, gives
$(e_{H^{\prime}}(X_{j}^{\prime},\overline{Z_{j}^{\prime}})-e_{H^{\prime}}(Y_{j}^{\prime},\overline{Z_{j}^{\prime}}))+2\left(e_{H^{\prime}}(X_{j}^{\prime})-e_{H^{\prime}}(Y_{j}^{\prime})\right)-d\left(|X_{j}^{\prime}|-|Y_{j}^{\prime}|\right)$
$=$ $\left(e_{H}(X_{j},\overline{Z_{j}}\right)-e_{H}(Y_{j},\overline{Z_{j}}))+2\left(e_{H}(X_{j})-e_{H}(Y_{j})\right)-d\left(|X_{j}|-|Y_{j}|\right)=0.$
It is easy to see that for $\ell\in[r]-\{i,j\}$, we have $e_{H^{\prime}}(X_{\ell}^{\prime},\overline{Z_{\ell}^{\prime}})=e_{H}(X_{\ell},\overline{Z_{\ell}})$, $e_{H^{\prime}}(Y_{\ell}^{\prime},\overline{Z_{\ell}^{\prime}})=e_{H}(Y_{\ell},\overline{Z_{\ell}})$, $e_{H^{\prime}}(X_{\ell}^{\prime})=e_{H}(X_{\ell})$, $e_{H^{\prime}}(Y_{\ell}^{\prime})=e_{H}(Y_{\ell})$, $|X_{\ell}|=|X_{\ell}^{\prime}|$, and $|Y_{\ell}|=|Y_{\ell}^{\prime}|$, proving (4) in this case.
Now consider the case when $(W_{i},W_{j})=(X_{i},X_{i})$. In this case, note that the following six equations hold.
$e_{H^{\prime}}(X_{i}^{\prime},\overline{Z_{i}^{\prime}})-e_{H}(X_{i},\overline{Z_{i}})$ $=-d_{H}(v,\overline{Z_{i}})$
$e_{H^{\prime}}(Y_{i}^{\prime},\overline{Z_{i}^{\prime}})-e_{H}(Y_{i},\overline{Z_{i}})$ $=d_{H}(v,\overline{Z_{i}})$
$e_{H^{\prime}}(X_{i}^{\prime})-e_{H}(X_{i})$ $=-d_{H}(v,X_{i})-(d-d_{H}(v))$
$e_{H^{\prime}}(Y_{i}^{\prime})-e_{H}(Y_{i})$ $=0$
$|X_{i}^{\prime}|-|X_{i}|$ $=-1$
$|Y_{i}^{\prime}|-|Y_{i}|$ $=1.$
As before, multiplying these equations by 1,1,2,2,d,d1,-1,2,-2,-$d,$ respectively, and adding up the resulting equations gives
$(e_{H^{\prime}}(X_{i}^{\prime},\overline{Z_{i}^{\prime}})-e_{H^{\prime}}(Y_{i}^{\prime},\overline{Z_{i}^{\prime}}))+2\left(e_{H^{\prime}}(X_{i}^{\prime})-e_{H^{\prime}}(Y_{i}^{\prime})\right)-d\left(|X_{i}^{\prime}|-|Y_{i}^{\prime}|\right)$
$=$ $\left(e_{H}(X_{i},\overline{Z_{i}}\right)-e_{H}(Y_{i},\overline{Z_{i}}))+2\left(e_{H}(X_{i})-e_{H}(Y_{i})\right)-d\left(|X_{i}|-|Y_{i}|\right)=0.$
Again, it is easy to see that for $\ell\in[r]-\{i\}$, we trivially have $e_{H^{\prime}}(X_{\ell}^{\prime},\overline{Z_{\ell}^{\prime}})=e_{H}(X_{\ell},\overline{Z_{\ell}})$, $e_{H^{\prime}}(Y_{\ell}^{\prime},\overline{Z_{\ell}^{\prime}})=e_{H}(Y_{\ell},\overline{Z_{\ell}})$, $e_{H^{\prime}}(X_{\ell}^{\prime})=e_{H}(X_{\ell})$, $e_{H^{\prime}}(Y_{\ell}^{\prime})=e_{H}(Y_{\ell})$, $|X_{\ell}|=|X_{\ell}^{\prime}|$, and $|Y_{\ell}|=|Y_{\ell}^{\prime}|$, proving (4) in this case and completing the proof of the claim. ∎
Let us bound the number of steps taken by the procedure Move-high-degree-vertices by the time it ends, that is, by the time (3) is no longer satisfied by any choice of parameters. Notice that after each step of the procedure, the number of edges in $H$ decreases by at least $\rho n$. Since $H$ satisfies (H4) at the beginning of the procedure, this implies that the total number of steps taken before the procedure ends is at most $2r\beta n^{2}/(\rho n)=(2r\beta/\rho)n\leq\rho n/2$ (using $r\leq\lceil 2/c\rceil$ and $\beta\ll\rho,\delta,c$). Since at most one vertex is moved at each step of the procedure we obtain the following.
  1. (M)
    At most $(2r\beta/\rho)n\leq\rho n/2$ vertices are moved during the procedure Move-high-degree-vertices.
In the rest of the proof, let $(X_{i})_{i\in[r]}$, $(Y_{i})_{i\in[r]}$, $(Z_{i})_{i\in[r]}$ and $H$ denote the parts and the graph obtained at the end of the procedure Move-high-degree-vertices. Then we claim that the following properties hold.
  1. (H5)
    $\delta(G[X_{i},Y_{i}])\geq\rho n/2$ for $i\in[r]$.
  2. (H6)
    For every $i,j\in[r]$, $W_{i}\in\{X_{i},Y_{i}\}$, $W_{j}\in\{X_{j},Y_{j}\}$ either $e_{H}(W_{i},W_{j})<d$ or $\Delta(H[W_{i},W_{j}])\leq\rho n$.
Indeed, note that for every vertex $v$ in $X_{i}\cup Y_{i}$, there is a step during the procedure Move-high-degree-vertices where its degree is at least $\rho n$ in $G[X^{*}_{i},Y^{*}_{i}]$, where $X^{*}_{i}$ and $Y^{*}_{i}$ denote the parts corresponding to $X_{i}$ and $Y_{i}$, respectively, in that step. (Here we used that if $v$ is moved into $X^{*}_{i}$ or $Y^{*}_{i}$ at the previous step, then it has degree at least $\rho n$ in $G[X^{*}_{i},Y^{*}_{i}]$, and if $v$ has never been moved during the procedure, then we use that (G2) holds at the beginning of the procedure, and that $\rho\ll\delta$.) Moreover, by (M), at most $\rho n/2$ vertices are moved during the procedure. Hence, the degree of $v$ changes by at most $\rho n/2$ when the procedure ends, proving that (H5) holds (note that (H5) pertains to the degree in $G$, not in $H$). Moreover, it is easy to see that (H6) holds, because otherwise the procedure would not have stopped.
In the next subsection, we carefully construct a small collection of vertex-disjoint copies of $K_{t,t}$ whose removal makes the subgraphs $G[X_{i},Y_{i}]$, for $i\in[r]$, more balanced.
4.2 Constructing a collection $\mathcal{K}$ of $K_{t,t}$’s in $G$ for balancing the subgraphs $G[X_{i},Y_{i}]$
In the rest of this proof, let
$T\coloneqq\frac{4t}{c}\left(\frac{4e}{\rho}\right)^{t}.$
By repeatedly applying an unbalanced variant of the Kővári–Sós–Turán theorem (Theorem 3.8), we prove the following claim. Recall that the vertex set of a given copy $K$ of $K_{t,t}$ in $G$ is denoted by $V(K)$, and the set of vertices contained in a given collection $\mathcal{K}$ of copies of $K_{t,t}$ in $G$ is denoted by $V(\mathcal{K})$.
Claim 4.3.
There exists a collection $\mathcal{K}\coloneqq\cup_{1\leq i\leq j\leq r}\big(\mathcal{K}(X_{i},X_{j})\cup\mathcal{K}(X_{i},Y_{j})\cup\mathcal{K}(Y_{i},X_{j})\cup\mathcal{K}(Y_{i},Y_{j})\big)$ of pairwise vertex-disjoint copies of $K_{t,t}$ in $G$, such that for every $i,j\in[r]$ with $i\leq j$ and every $W_{i}\in\{X_{i},Y_{i}\}$, $W_{j}\in\{X_{j},Y_{j}\}$, the following two properties hold.
  • If $e_{H}(W_{i},W_{j})\geq Td$ then $\big|\mathcal{K}(W_{i},W_{j})\big|=\left\lfloor\frac{e_{H}(W_{i},W_{j})}{d}\right\rfloor$ and, otherwise, $\mathcal{K}(W_{i},W_{j})=\emptyset$. In particular, $|\mathcal{K}|\leq\sum_{i\leq j}\frac{e_{H}(W_{i},W_{j})}{d}\leq\frac{e(H)}{d}$.
  • Every copy $K\in\mathcal{K}(W_{i},W_{j})$ satisfies $V(K)\subseteq Z_{i}\cup Z_{j}$, and for every $\ell\in\{i,j\}$, we have
    $|W_{\ell}\cap V(K)|-|U_{\ell}\cap V(K)|=\left\{\begin{array}[]{ll}1&\text{if $i\neq j$}\\ 2&\text{if $i=j$,}\end{array}\right$. (5)
    where $U_{\ell}=Z_{\ell}-W_{\ell}$.
Let $\sigma$ be an ordering of the pairs $(W_{i},W_{j})$, with $1\leq i\leq j\leq r$ and $e_{H}(W_{i},W_{j})\geq Td$, in increasing order of $e_{H}(W_{i},W_{j})$. We will show by induction that for every such pair $(W_{i},W_{j})$ there is a collection $\mathcal{K}$ of pairwise vertex-disjoint copies of $K_{t,t}$ in $G$ satisfying the following: for every $(W_{i^{\prime}},W_{j^{\prime}})$ such that $(W_{i^{\prime}},W_{j^{\prime}})\leq_{\sigma}(W_{i},W_{j})$, $\mathcal{K}$ consists of exactly $\left\lfloor\frac{e_{H}(W_{i^{\prime}},W_{j^{\prime}})}{d}\right\rfloor$ copies $K$ of $K_{t,t}$ satisfying $V(K)\subseteq Z_{i^{\prime}}\cup Z_{j^{\prime}}$ and (5).
To this end, fix $(W_{i},W_{j})$ such that $1\leq i\leq j\leq r$ and $e_{H}(W_{i},W_{j})\geq Td$, and suppose that there is a collection $\mathcal{K}$ of pairwise vertex-disjoint copies of $K_{t,t}$ in $G$ satisfying the following: for every $(W_{i^{\prime}},W_{j^{\prime}})$ such that $(W_{i^{\prime}},W_{j^{\prime}})<_{\sigma}(W_{i},W_{j})$, $\mathcal{K}$ consists of exactly $\left\lfloor\frac{e_{H}(W_{i^{\prime}},W_{j^{\prime}})}{d}\right\rfloor$ copies $K$ of $K_{t,t}$ satisfying $V(K)\subseteq Z_{i^{\prime}}\cup Z_{j^{\prime}}$ and (5) (and there are no other $K_{t,t}$-copies). Let $f\coloneqq e_{H}(W_{i},W_{j})$, and let $\mathcal{K}^{\prime}$ be a maximal collection of at most $\left\lfloor f/d\right\rfloor$ copies $K$ of $K_{t,t}$ that are pairwise vertex-disjoint and vertex-disjoint from the copies of $K_{t,t}$ in $\mathcal{K}$ and satisfy $V(K)\subseteq Z_{i}\cup Z_{j}$ and (5).
If $|\mathcal{K}^{\prime}|=\left\lfloor f/d\right\rfloor$, then we are done with the proof of the induction step, so suppose $|\mathcal{K}^{\prime}|<\left\lfloor f/d\right\rfloor$. Let $V$ be the set of all vertices contained in a copy of $K_{t,t}$ in $\mathcal{K}\cup\mathcal{K}^{\prime}$. Then $|V|\leq(2r)^{2}\cdot(f/d)\cdot 2t=8r^{2}ft/d$ because each pair $(W_{i^{\prime}},W_{j^{\prime}})$ that precedes $(W_{i},W_{j})$ in $\sigma$ satisfies $e_{H}(W_{i^{\prime}},W_{j^{\prime}})\leq f$ by the choice of $\sigma$, and there are at most $(2r)^{2}$ such pairs $(W_{i^{\prime}},W_{j^{\prime}})$. Thus, the number of edges in $H[W_{i},W_{j}]$ incident to a vertex from $V$ is at most $\rho n\cdot|V|\leq\rho n\cdot 8r^{2}ft/d\leq f/2$ (using that $\Delta(H[W_{i},W_{j}])\leq\rho n$ by (H6), that $r\leq\left\lceil 1/c\right\rceil$ and that $\rho\ll c,1/t$). This implies that there is a vertex $v$ in $W_{i}$ with at least $\frac{f}{2n}\geq\frac{Td}{2n}\geq 2t\left(\frac{4e}{\rho}\right)^{t}$ neighbours in $W_{j}-V$; denote the set of these neighbours by $A$. Observe that $G[W_{j}-V,U_{j}-V]$ has minimum degree at least $\rho n/4$ because (H5) holds (by Claim 4.2) and $|V|\leq 2t\cdot e(H)/d\leq 2t\cdot 2r\beta n^{2}/d\leq\rho n/4$ because (H4) holds, $d\geq cn$ and $\beta\ll\rho,c,1/t$. Thus, by the Kővári–Sós–Turán theorem (Lemma 3.8), there is a copy of $K_{t,t-1}$ in $G[A,U_{j}-V]$, with $t$ vertices in $A$ and t1t-1 vertices in $U_{j}-V$. Together with $v$, this copy of $K_{t,t-1}$ forms a copy $K$ of $K_{t,t}$ in $G$ satisfying $V(K)\subseteq Z_{i}\cup Z_{j}$ and (5), contradicting the maximality of the collection $\mathcal{K}^{\prime}$, and proving the claim. ∎
Let $\mathcal{K}$ be the collection of copies of $K_{t,t}$ guaranteed by Claim 4.3. We claim that, for each $i\in[r]$, we have
$|X_{i}\cap V(\mathcal{K})|-|Y_{i}\cap V(\mathcal{K})|=|X_{i}|-|Y_{i}|\pm 8rT.$ (6)
Indeed, let us fix $i\in[r]$, and note that Claim 4.3 ensures that for every $j\in[r]$, $W_{i}\in\{X_{i},Y_{i}\}$, $W_{j}\in\{X_{j},Y_{j}\}$ such that $e_{H}(W_{i},W_{j})\geq Td$ and $W_{j}\neq Z_{i}-W_{i}$, there are exactly $\left\lfloor\frac{e_{H}(W_{i},W_{j})}{d}\right\rfloor$ copies $K\in\mathcal{K}$ satisfying $V(K)\subseteq Z_{i}\cup Z_{j}$ and (5). Therefore, it is easy to see that the following holds (where $\mathbbm{1}\{E\}$ denotes the indicator function of $E$ i.e., $\mathbbm{1}\{E\}=1$ if $E$ holds, and $\mathbbm{1}\{E\}=0$, otherwise).
$|X_{i}\cap V(\mathcal{K})|-|Y_{i}\cap V(\mathcal{K})|$
$\quad=\,\sum_{W_{j}\in\{X_{j},Y_{j}\},\,\,j\in[r]-\{i\}}\left(\left\lfloor\frac{e_{H}(X_{i},W_{j})}{d}\right\rfloor\mathbbm{1}\{e_{H}(X_{i},W_{j})\geq Td\}-\left\lfloor\frac{e_{H}(Y_{i},W_{j})}{d}\right\rfloor\mathbbm{1}\{e_{H}(Y_{i},W_{j})\geq Td\}\right)$
$\qquad\quad+\left(2\left\lfloor\frac{e_{H}(X_{i})}{d}\right\rfloor\mathbbm{1}\{e_{H}(X_{i})\geq Td\}-2\left\lfloor\frac{e_{H}(Y_{i})}{d}\right\rfloor\mathbbm{1}\{e_{H}(Y_{i})\geq Td\}\right)$
$\quad=\,\sum_{W_{j}\in\{X_{j},Y_{j}\},\,\,j\in[r]-\{i\}}\left(\frac{e_{H}(X_{i},W_{j})}{d}-\frac{e_{H}(Y_{i},W_{j})}{d}\right)+\left(\frac{2e_{H}(X_{i})}{d}-\frac{2e_{H}(Y_{i})}{d}\right)\pm 8rT$
$\quad=\,\frac{1}{d}\cdot\left(e_{H}(X_{i},\overline{Z_{i}})-e_{H}(Y_{i},\overline{Z_{i}})+2e_{H}(X_{i})-2e_{H}(Y_{i})\right)\pm 8rT=|X_{i}|-|Y_{i}|\pm 8rT,$
where for the last equality we used that (H1) holds (by Claim 4.2). Hence (6) holds.
Note that by (6), for every $i\in[r]$, we have
$|X_{i}-V(\mathcal{K})|=|Y_{i}-V(\mathcal{K})|\pm 8rT.$
For every $i\in[r]$, let $X_{i}^{\prime}\subseteq X_{i}-V(\mathcal{K})$ and $Y_{i}^{\prime}\subseteq Y_{i}-V(\mathcal{K})$ be subsets of equal size, obtained by removing at most $8rT$ vertices from $X_{i}-V(\mathcal{K})$ and from $Y_{i}-V(\mathcal{K})$. Write $Z_{i}^{\prime}=X_{i}^{\prime}\cup Y_{i}^{\prime}$. Then, since $\{Z_{1},\ldots,Z_{r}\}$ partition $V(G)$, we have
$|V(G)-(Z_{1}^{\prime}\cup\ldots\cup Z_{r}^{\prime})|\leq|V(\mathcal{K})|+8rT\cdot 2r=|V(\mathcal{K})|+64r^{2}\cdot\frac{t}{c}\cdot\left(\frac{4e}{\rho}\right)^{t}.$ (7)
Let $G^{\prime}$ be the spanning subgraph of $G$ obtained by removing the edges of $G$ within $X_{i}^{\prime}$ and $Y_{i}^{\prime}$ for all $i\in[r]$ such that $G[Z^{\textrm{start}}_{i}]$ is $\beta$-almost-bipartite. This ensures that the subgraph $G^{\prime}[Z_{i}^{\prime}]$ is bipartite and balanced for all $i\in[r]$ for which $G[Z^{\textrm{start}}_{i}]$ is $\beta$-almost-bipartite. (Recall that here the sets $Z^{\textrm{start}}_{i}$ for $i\in[r]$, are defined before running the procedure Move-high-degree-vertices.)
Let us make a simple observation concerning the number of vertices added to or removed from the set $Z_{i}^{\rm{start}}$ in order to form the set $Z_{i}^{\prime}$ for every $i\in[r]$. Note that, for every $i\in[r]$, by (M) and the definition of $Z_{i}^{\prime}$, we have $|Z_{i}^{\rm{start}}-Z_{i}^{\prime}|\leq(2r\beta/\rho)n+|V(\mathcal{K})|+16rT$ and $|Z_{i}^{\prime}-Z_{i}^{\rm{start}}|\leq(2r\beta/\rho)n$. Moreover, since $r\leq\left\lceil 1/c\right\rceil$ and $\beta\ll\rho,\sigma,c,\delta$, we have $(2r\beta/\rho)n\leq\sigma n/4$, and by Claim 4.3, (H4) and the fact that $\beta\ll\rho,\sigma,c,1/t$, we have $|V(\mathcal{K})|\leq 2t\cdot\frac{e(H)}{d}\leq 2t\cdot\frac{2r\beta n^{2}}{d}\leq\sigma n/4$. Therefore,
$|Z_{i}^{\rm{start}}-Z_{i}^{\prime}|,|Z_{i}^{\prime}-Z_{i}^{\rm{start}}|\leq\sigma n.$ (8)
4.3 Showing that expansion is preserved after balancing
In this subsection, we will prove the following claim, which shows that the subgraphs $G^{\prime}[Z_{i}^{\prime}]$ for $i\in[r]$ (which are obtained after balancing) still have good expansion properties.
Claim 4.4.
For every $i\in[r]$, the graph $G^{\prime}[Z_{i}^{\prime}]$ has no $\rho/8$-sparse cuts.
Fix $i\in[r]$, and, for convenience, write $Z^{\prime}\coloneqq Z_{i}^{\textrm{start}}$ and $Z\coloneqq Z_{i}^{\prime}$. Consider a partition $\{A,B\}$ of $Z$, where $A$ and $B$ are non-empty and $|A|\geq|B|$. We will show that $e_{G^{\prime}}(A,B)\geq(\rho/8)|A||B|$, thereby proving the claim.
Suppose first that $|B|\leq\rho n/8$. Recall that $G[X_{i},Y_{i}]$ has minimum degree at least $\rho n/2$ (by (H5)). Since $|X_{i}-X_{i}^{\prime}|,|Y_{i}-Y_{i}^{\prime}|\leq\sigma n\leq\rho n/4$, by (8) and $\sigma\ll\rho$, the graph $G[X_{i}^{\prime},Y_{i}^{\prime}]$ has minimum degree at least $\rho n/4$. As this is a spanning subgraph of $G^{\prime}[Z_{i}^{\prime}]=G^{\prime}[Z]$, the graph $G^{\prime}[Z]$ also has minimum degree at least $\rho n/4$. Hence
$e_{G^{\prime}}(A,B)\geq|B|\cdot\big(\rho n/4-|B|\big)\geq|B|\cdot\rho n/8\geq(\rho/8)\cdot|A||B|,$
Suppose, instead, that $|B|>\rho n/8$. Write $U\coloneqq Z-Z^{\prime}$, $W\coloneqq Z^{\prime}-Z$, $A^{\prime}\coloneqq(A-U)\cup W$ and $B^{\prime}\coloneqq B-U$. Then $\{A^{\prime},B^{\prime}\}$ is a partition of $Z^{\prime}$. Since $G[Z^{\prime}]$ has no $\zeta$-sparse cuts, we have
$\begin{split}e_{G}(A^{\prime},B^{\prime})\geq\zeta|A^{\prime}||B^{\prime}|\geq\zeta\cdot|A-U|\cdot|B-U|\geq\zeta\cdot(|A|-\sigma n)\cdot(|B|-\sigma n)\geq(\zeta/4)\cdot|A||B|.\end{split}$ (9)
For the last inequality, we used $|A|\geq|B|\geq\rho n/8\geq 2\sigma n$ (using $\sigma\ll\rho$). We now upper bound the number of edges in $G[Z^{\prime}]$ that are not in $G^{\prime}[Z^{\prime}]$. Write $X^{\prime}\coloneqq X_{i}^{\mathrm{start}}$, $Y^{\prime}\coloneqq Y_{i}^{\mathrm{start}}$, $X\coloneqq X_{i}$ and $Y\coloneqq Y_{i}$. Recall that if $G^{\prime}[Z]$ and $G[Z]$ differ, then $e_{G}(X^{\prime})+e_{G}(Y^{\prime})\leq\beta n^{2}$ and $G^{\prime}[Z]$ is obtained by removing some of the edges in $G[X]$ and $G[Y]$ from $G[Z]$. Assuming $G[Z]$ and $G^{\prime}[Z]$ differ, we thus have
$\begin{split}e_{G}(Z)-e_{G^{\prime}}(Z)=e_{G}(X)+e_{G}(Y)&\leq e_{G}(X^{\prime})+|X-X^{\prime}|\cdot n+e_{G}(Y^{\prime})+|Y-Y^{\prime}|\cdot n\\ &\leq\beta n^{2}+2\sigma n^{2}\leq 3\sigma n^{2},\end{split}$ (10)
using (8) and $\beta\ll\sigma$. Hence
$e_{G^{\prime}}(A,B)=e_{G^{\prime}}(A\cup W,B)-e_{G^{\prime}}(W,B)$ $\geq e_{G^{\prime}}(A^{\prime},B^{\prime})-\sigma n^{2}$
$\geq e_{G}(A^{\prime},B^{\prime})-4\sigma n^{2}$
$\geq(\zeta/4)\cdot|A||B|-4\sigma n^{2}\geq(\zeta/8)\cdot|A||B|\geq(\rho/8)\cdot|A||B|.$
Here we used $|W|\leq\sigma n$ for the first inequality, (10) for the second one, (9) for the third, $(\zeta/4)\cdot|A||B|\geq(\zeta\rho^{2}/256)\cdot n^{2}\geq 8\sigma n^{2}$ for the fourth, and that $\rho\ll\zeta$ for the last inequality. ∎
4.4 Putting everything together and completing the proof of Lemma 4.1
To complete the proof of Lemma 4.1, we will show that the following properties hold for all $i\in[r]$.
  1. (a)
    $G^{\prime}[Z_{i}^{\prime}]$ has no $\rho/8$-sparse cuts.
  2. (b)
    Either $G^{\prime}[Z_{i}^{\prime}]$ is bipartite and balanced, or one needs to remove at least $\gamma n^{2}/4$ edges from $G^{\prime}[Z_{i}^{\prime}]$ to make it bipartite.
  3. (c)
    The average degree of $G^{\prime}[Z_{i}^{\prime}]$ is at least $d-\frac{144\sigma}{\rho^{2}}|Z_{i}^{\prime}|$.
Indeed, Claim 4.4 shows (a).
For proving (b), notice that if $G[Z_{i}^{\textrm{start}}]$ is $\beta$-almost-bipartite, then $G^{\prime}[Z_{i}^{\prime}]$ is indeed bipartite and balanced, by the choice of $X_{i}^{\prime}$ and $Y_{i}^{\prime}$. So consider $i\in[r]$ such that $G[Z_{i}^{\textrm{start}}]$ is $\gamma$-far-from-bipartite. Recall that $|Z_{i}^{\textrm{start}}-Z_{i}^{\prime}|\leq\sigma n\leq\gamma n/4$ by (8) and the fact that $\sigma\ll\gamma$. Suppose for a contradiction that one can remove fewer than $\gamma n^{2}/4$ edges from $G^{\prime}[Z_{i}^{\prime}]=G[Z_{i}^{\prime}]$ to make it bipartite. But this means one can remove fewer than $\gamma n^{2}/4+|Z_{i}^{\textrm{start}}-Z_{i}^{\prime}|\cdot n\leq\gamma n^{2}/2$ edges from $G[Z_{i}^{\textrm{start}}]$ to make it bipartite. (Indeed, we can remove all edges incident to $Z_{i}^{\textrm{start}}-Z_{i}^{\prime}$, and fewer than $\gamma n^{2}/4$ edges from $G[Z_{i}^{\textrm{start}}\cap Z_{i}^{\prime}]$ to make it bipartite.) This contradicts the fact that $G[Z_{i}^{\textrm{start}}]$ is $\gamma$-far-from-bipartite, and proves (b).
Finally, for proving (c), first recall that $e_{G}(Z_{i}^{\prime})-e_{G^{\prime}}(Z_{i}^{\prime})\leq 3\sigma n^{2}$, by (10). Second, by (8), and since the sets $Z_{i}^{\textrm{start}}$ satisfy (E1), we have
$e_{G}(Z_{i}^{\prime},\overline{Z_{i}^{\prime}})\leq e_{G}(Z_{i}^{\textrm{start}},\overline{Z_{i}^{\textrm{start}}})+|Z_{i}^{\textrm{start}}-Z_{i}^{\prime}|n+|Z_{i}^{\prime}-Z_{i}^{\textrm{start}}|n\leq\eta n^{2}+2\sigma n^{2}\leq 3\sigma n^{2}.$
Combining the above two inequalities we have
$d|Z_{i}^{\prime}|-2e_{G^{\prime}}(Z_{i}^{\prime})\leq e_{G}(Z_{i}^{\prime},\overline{Z_{i}^{\prime}})+2(e_{G}(Z_{i}^{\prime})-e_{G^{\prime}}(Z_{i}^{\prime}))\leq 9\sigma n^{2}\leq\frac{144\sigma}{\rho^{2}}|Z_{i}^{\prime}|^{2},$ (11)
where in the last inequality we used $|Z_{i}^{\prime}|\geq\rho n/4$, which follows from $G[Z_{i}^{\prime}]$ having minimum degree at least $\rho n/4$, as pointed out at the beginning of the proof of Claim 4.4. Hence, by (11), the average degree of $G^{\prime}[Z_{i}^{\prime}]$ is at least $d-\frac{144\sigma}{\rho^{2}}|Z_{i}^{\prime}|$, proving (c).
Using (a), (b) and (c), it is now easy to check that (B2), (B3) and (B4) of the lemma hold with $Z_{i}^{\prime}$ playing the role of $Z_{i}$ for $i\in[r]$, $V(\mathcal{K})$ playing the role of $L$ and with $\frac{144\sigma}{\rho^{2}},\rho/8$ and $\gamma/4$ playing the roles of $\eta,\zeta$ and $\gamma$, respectively (in particular, notice that these new parameters satisfy $\eta\ll\zeta\ll\gamma\ll c$, as $\sigma\ll\rho\ll\gamma\ll c$). Since, by definition, $\mathcal{K}$ is a collection of pairwise vertex-disjoint copies of $K_{t,t}$ in $G$, (B5) of the lemma immediately follows. Finally, by (7) and using that $\rho/8$ plays the role of $\zeta$, we have
$\big|V(G)-\big(Z_{1}^{\prime}\cup\ldots\cup Z_{r}^{\prime}\cup V(\mathcal{K})\big)\big|\leq 64r^{2}\cdot\frac{t}{c}\cdot\left(\frac{e}{2\zeta}\right)^{t},$
proving (B1), and completing the proof of Lemma 4.1. ∎
5 Packing $K_{t,t}$’s in expanders
Our second key lemma for proving Theorem 1.4 is as follows. This lemma shows that expanders admit a perfect $K_{t,t}$-packing if their number of vertices is divisible by $2t$. It will be used to find a $K_{t,t}$-packing that covers all but at most 2t12t-1 vertices in each of the expanders provided by Lemma 4.1.
Lemma 5.1.
Let $t\geq 2$, $n_{0}$ be integers, let $0<c,\gamma,\zeta,\eta<1$ and suppose that $1/n_{0}\ll\eta\ll\zeta\ll\gamma\ll c,1/t$. Let $n\geq n_{0}$ be an integer divisible by $2t$, let $d\geq cn$ and let $G$ be an $n$-vertex graph with the following properties.
  1. (Q1)
    $G$ has average degree at least $d-\eta n$ and maximum degree at most $d$.
  2. (Q2)
    $G$ has no $\zeta$-sparse cuts.
  3. (Q3)
    $G$ is either bipartite and balanced, or $\gamma$-far-from-bipartite.
Then $G$ has a perfect $K_{t,t}$-packing.
Note that the following corollary, where we drop the assumption of divisibility from Lemma 5.1 and allow for up to 2t12t-1 uncovered vertices, follows directly from Lemma 5.1. Indeed, given $G$ as in the next corollary, remove up to 2t12t-1 vertices to obtain a graph $G^{\prime}$ whose number of vertices $n^{\prime}$ is divisible by $2t$. It is easy to see that the properties (Q1)(Q3) above hold for $G^{\prime}$, with slightly worse parameters, say, with $n_{0}-2t,c/2,\gamma/2,2\eta,\zeta/2,n^{\prime},d$ in place of $n_{0},c,\gamma,\eta,\zeta,n,d$ respectively.
Corollary 5.2.
Let $t\geq 2$, $n_{0}$ be integers, let $0<c,\gamma,\zeta,\eta<1$ and suppose that $1/n_{0}\ll\eta\ll\zeta\ll\gamma\ll c,1/t$. Let $n\geq n_{0}$, let $d\geq cn$ and let $G$ be an $n$-vertex graph with the following properties.
  1. (1)
    $G$ has average degree at least $d-\eta n$ and maximum degree at most $d$.
  2. (2)
    $G$ has no $\zeta$-sparse cuts.
  3. (3)
    $G$ is either bipartite and balanced, or $\gamma$-far-from-bipartite.
Then, $G$ has a $K_{t,t}$-packing covering all but at most 2t12t-1 vertices.
Let us now proceed with the proof of Lemma 5.1.
Let $\varepsilon,\mu,\xi$ be positive numbers satisfying
$1/n_{0}\ll\varepsilon\ll\mu\ll\eta\ll\xi\ll\zeta.$
Apply the regularity lemma (Lemma 3.5) to $G$ with parameters $\varepsilon$ and $\mu$, to obtain positive numbers $M=M(\varepsilon)$, $m$ with $m\leq M$, a subgraph $G^{\prime}$ and a partition $\{V_{0},\ldots,V_{m}\}$ of $V(G)$ into clusters $V_{1},\ldots,V_{m}$ and an exceptional set $V_{0}$ satisfying properties (R1)(R6). Let $\Gamma$ be the corresponding reduced graph, namely, it has vertex set $\{V_{1},\ldots,V_{m}\}$ and $V_{i}V_{j}$ is an edge in $\Gamma$ if and only if $G^{\prime}[V_{i},V_{j}]$ is $\varepsilon$-regular with density more than $\mu$. Equivalently, $V_{i}V_{j}$ is an edge in $\Gamma$ if and only if there is an edge of $G^{\prime}$ between $V_{i}$ and $V_{j}$ (by (R6)).
Notice that if $G$ is bipartite, then both $G^{\prime}$ and $\Gamma$ are bipartite. Indeed, if $G$ is bipartite then $G^{\prime}$ is bipartite since it is a subgraph of $G$, and $\Gamma$ is bipartite because every $V_{i}$ with $i\in[m]$ is contained in one of the parts of $G$ by (R2).
In the rest of the proof, we will assume that $m$ is even, and that if $G$ is bipartite and balanced, then both $G^{\prime}$ and $\Gamma$ are bipartite and balanced. Indeed, since all clusters $V_{i}$ with $i\in[m]$ are equal in size — each of size at most $\varepsilon n$ by (R3) — and each $V_{i}$ lies entirely within one of the two parts of the bipartition of $G$ by (R2), this adjustment can be made, if necessary, by moving the vertices of some clusters, covering up to $\varepsilon n$ vertices, from $G^{\prime}$ into $V_{0}$. Moreover, after (potentially) moving these vertices, it still follows from (R3) that
$|V_{0}|\leq 2\varepsilon n,$ (12)
and by (R4), for every $v\in V(G^{\prime})$ we have
$d_{G^{\prime}}(v)\geq d_{G}(v)-(\mu+2\varepsilon)n.$ (13)
We claim that, after this potential adjustment, $G^{\prime}$ satisfies the following properties.
  1. (G’1)
    $G^{\prime}$ has average degree at least $d-3\eta n$.
  2. (G’2)
    $G^{\prime}$ has no $\zeta/4$-sparse cuts. (In particular, $G^{\prime}$ has minimum degree at least $\zeta/4\cdot(|V(G^{\prime})|-1)\geq\zeta n/5.$)
  3. (G’3)
    If $G$ is bipartite and balanced, then $G^{\prime}$ is bipartite and balanced. Otherwise, one needs to remove at least $\gamma n^{2}/2$ edges from $G^{\prime}$ to make it bipartite. (In particular, $G^{\prime}$ is $\gamma/2$-far-from-bipartite.)
Indeed, (G’1) follows directly from (13) and the fact that the average degree of $G$ is at least $d-\eta n$ and $\varepsilon,\mu\ll\eta$. To prove (G’2), we use (13) together with the fact that $G$ has no $\zeta$-sparse cuts, as follows. For every partition $\{X,Y\}$ of $V(G^{\prime})$ with $|X|\leq|Y|$, we have
$e_{G^{\prime}}(X,Y)$ $\geq e_{G}(X,V(G)-X)-|X|(\mu+2\varepsilon)n$
$\geq\zeta|X|\cdot(n-|X|)-|X|(\mu+2\varepsilon)n$
$\geq\left(\frac{\zeta}{2}-(\mu+2\varepsilon)\right)\cdot|X|\cdot n\geq\frac{\zeta}{4}\cdot|X|\cdot|Y|,$
using $|X|\leq n/2$ for the third inequality, and $\varepsilon,\mu\ll\zeta$ for the last inequality. This proves (G’2). Finally, to prove (G’3), observe that if $G$ is bipartite and balanced, then, as noted before, we can assume that $G^{\prime}$ is also bipartite and balanced. Otherwise, by our assumption in Lemma 5.1, $G$ is $\gamma$-far-from-bipartite, in which case we claim that at least $\gamma n^{2}/2$ edges need to be removed from $G^{\prime}$ to make it bipartite. Indeed, by (12) and (13), the number of edges in $G$ that are not present in $G^{\prime}$ is at most $|V_{0}|\cdot n+n\cdot(\mu+2\varepsilon)n\leq(\mu+4\varepsilon)n^{2}\leq\gamma n^{2}/2$. Since $G$ is $\gamma$-far-from-bipartite, at least $\gamma n^{2}$ edges must be removed to make it bipartite. Therefore, even after accounting for the fact that at most $\gamma n^{2}/2$ of these edges are missing in $G^{\prime}$, one still needs to remove at least $\gamma n^{2}/2$ edges from $G^{\prime}$, proving (G’3).
5.1 Finding vertex-disjoint super-regular subgraphs that cover all vertices outside $V_{0}$
Recall that the clusters $V_{i}$, $i\in[m]$, have the same size by (R3). For each $i\in[m]$, if $|V_{i}|$ is odd, move one vertex from $V_{i}$ to $V_{0}$. Now for each $i\in[m]$, partition $V_{i}$ into two sets $V_{i}^{\prime},V_{i}^{\prime\prime}$ of equal size, let $\Gamma_{\rm{ref}}$ be the graph with the vertex set $\{V_{1}^{\prime},V_{1}^{\prime\prime},\ldots,V_{m}^{\prime},V_{m}^{\prime\prime}\}$ whose edges are all the pairs $AB$ with $A\in\{V_{i}^{\prime},V_{i}^{\prime\prime}\}$, $B\in\{V_{j}^{\prime},V_{j}^{\prime\prime}\}$ where $V_{i}V_{j}$ is an edge in $\Gamma$. (In particular, $\Gamma_{\rm{ref}}$ is the $2$-lift of $\Gamma$.) Then note that every edge in $\Gamma_{\rm{ref}}$ corresponds to a $3\varepsilon$-regular subgraph of $G^{\prime}$ with a density more than $\mu-\varepsilon\geq\mu/2$.
We claim that $\Gamma_{\rm{ref}}$ contains a perfect matching. To see this, we first prove the following claim. A $2$-matching is a collection of pairwise vertex-disjoint edges and odd cycles. A perfect $2$-matching in a graph is a $2$-matching covering all vertices in the graph.
Claim 5.3.
$\Gamma$ has a perfect $2$-matching.
Before we prove the claim, recall that a fractional matching of a graph $G$ is an edge weighting $w:E(G)\to[0,1]$, such that $\sum_{u\in N_{G}(v)}w(uv)\leq 1$, for every vertex $v$ in $G$. A perfect fractional matching is a fractional matching where we have equality $\sum_{u\in N_{G}(v)}w(uv)=1$ for every vertex $v$ in $G$. We need the following useful fact.
Fact 5.4.
Let $G$ be a graph which has a perfect fractional matching. Then it has a perfect $2$-matching.
This is a standard fact, but for completeness we prove it in Appendix A.
By (G’1)(G’3) and Lemma 3.3, the graph $G^{\prime}$ has a Hamilton cycle, which we denote by $C=(v_{1}\ldots v_{n^{\prime}})$, where $n^{\prime}=|V(G^{\prime})|$. Let $w:E(\Gamma)\to\mathbb{R}^{\geq 0}$ be the edge-weighting where $w(V_{i}V_{j})$ is the number of edges in $C$ with one end in $V_{i}$ and the other in $V_{j}$, divided by $2|V_{1}|$. We claim that $w$ is a perfect fractional matching in $\Gamma$; that is, $\sum_{j\in[m]}w(V_{i}V_{j})=1$ for every $i\in[m]$. Indeed, for every $i\in[m]$, since $G^{\prime}$ has no edges with both ends in $V_{i}$, the number of edges in $C$ with an end in $V_{i}$ is exactly $2|V_{i}|=2|V_{1}|$, so $\sum_{j\in[m]}w(V_{i}V_{j})=1$ for every $i\in[m]$, as claimed. By Fact 5.4, this implies that $\Gamma$ has a perfect $2$-matching, as required. ∎
Let $\mathcal{M}$ be a perfect $2$-matching in $\Gamma$. Replace each isolated edge $V_{i}V_{j}$ in $\mathcal{M}$ with two vertex-disjoint edges $V_{i}^{\prime}V_{j}^{\prime\prime}$ and $V_{i}^{\prime\prime}V_{j}^{\prime}$ in $\Gamma_{\rm{ref}}$ and each odd cycle $(V_{i_{1}},\ldots,V_{i_{2k+1}})$ (for some $k\geq 1$) in $\mathcal{M}$ with the edges $V_{i_{j}}^{\prime}V_{i_{j+1}}^{\prime\prime}\in E(\Gamma_{\rm{ref}})$ for $j\in[2k+1]$ (where addition in the indices is taken modulo $2k+1$), resulting in a perfect matching $\mathcal{M}_{\rm{ref}}$ in $\Gamma_{\rm{ref}}$.
Let $U_{1}U_{2},\ldots,U_{2m-1}U_{2m}$ denote the edges of $\mathcal{M}_{\rm{ref}}$. Since every edge in $\Gamma_{\rm{ref}}$ corresponds to a $3\varepsilon$-regular subgraph of $G^{\prime}$ with density more than $\mu/2$, it follows that for every $i\in[m]$, $G^{\prime}[U_{2i-1},U_{2i}]$ is $3\varepsilon$-regular with density more than $\mu/2$. Furthermore, by (R6), for all $i,j\in[2m]$, the graph $G^{\prime}[U_{i},U_{j}]$ is $3\varepsilon$-regular with density either 0 or more than $\mu/2$.
By Proposition 3.7, we can move $3\varepsilon|U_{i}|$ vertices from $U_{i}$ (for every $i\in[2m]$) into the set $V_{0}$, so that the resulting subgraphs $G^{\prime}[U_{2i-1},U_{2i}]$ are $(4\varepsilon,\mu/4)$-super-regular for every $i\in[m]$. By moving less than $2t$ additional vertices from each $U_{i}$ to $V_{0}$, we also assume that $|U_{i}|$ is divisible by $2t$ for every $i\in[2m]$. We claim that after moving these vertices, the following properties hold.
  1. (G’4)
    We have $\bigcup_{i=1}^{2m}U_{i}=V(G)-V_{0}=V(G^{\prime})$, $|U_{1}|=\ldots=|U_{2m}|$, and $|V_{0}|,|U_{1}|,\ldots,|U_{2m}|$ are divisible by $2t$. Furthermore, if $G$ is bipartite and balanced with the bipartition $\{X,Y\}$, then we can assume that $\bigcup_{i=1}^{m}U_{2i-1}=X-V_{0}$ and $\bigcup_{i=1}^{m}U_{2i}=Y-V_{0}$ (and hence $G^{\prime}$ is still bipartite and balanced).
  2. (G’5)
    For every $i\in[m]$, $G^{\prime}[U_{2i-1},U_{2i}]$ is $(5\varepsilon,\mu/5)$-super-regular.
  3. (G’6)
    For all $i,j\in[2m]$, the graph $G^{\prime}[U_{i},U_{j}]$ is $5\varepsilon$-regular with density either 0 or more than $\mu/5$.
Items (G’5) and (G’6) follow easily from the discussion above. To see why (G’4) holds, recall that, by (R3), the clusters $V_{i}$, $i\in[m]$ have the same size, and by definition, the sets $U_{i}$, $i\in[2m]$, had the same size before an equal number of vertices from each of these sets are moved to $V_{0}$. Furthermore, if $G$ is bipartite and balanced with the bipartition $\{X,Y\}$, then every cluster $V_{i}$, $i\in[m]$ is contained in one of the parts of $G$ by (R2), and moreover, as noted earlier, $G^{\prime}$ and the reduced graph $\Gamma$ are both bipartite and balanced. Hence, the $2$-matching $\mathcal{M}$ in $\Gamma$ is, in fact, a perfect matching in this case. Hence, $\mathcal{M}_{\rm ref}$ is also a perfect matching. Thus, without loss of generality, even after moving the vertices, we may assume $\bigcup_{i=1}^{m}U_{2i-1}=X-V_{0}$ and $\bigcup_{i=1}^{m}U_{2i}=Y-V_{0}$, as desired, proving (G’4).
Since a total of at most $m+3\varepsilon n+4tm\leq 4\varepsilon n$ vertices have been moved into $V_{0}$ from $G^{\prime}$, by (12), we have that
$|V_{0}|\leq 6\varepsilon n,$ (14)
and, by (G’1), we have that
  1. (G’7)
    $G^{\prime}$ has average degree at least $d-3\eta n-4\varepsilon n\geq d-6\eta n$.
Furthermore, we claim that the following two properties hold.
  1. (G’8)
    $G^{\prime}$ has no $\zeta/12$-sparse cuts. Moreover, $G^{\prime}$ has minimum degree at least $\zeta n/6.$
  2. (G’9)
    If $G$ is bipartite and balanced, then $G^{\prime}$ is bipartite and balanced. Otherwise, one needs to remove at least $\gamma n^{2}/4$ edges from $G^{\prime}$ to make it bipartite.
To see why (G’8) holds, first note that after moving at most $4\varepsilon n$ vertices from $G^{\prime}$ to $V_{0}$, the minimum degree of $G^{\prime}$ is at least $\zeta n/5-4\varepsilon n\geq\zeta n/6$ by (G’2). Consider a partition $\{X,Y\}$ of $G^{\prime}$ with $|X|\leq|Y|$ and note that if $|X|\leq\zeta n/12$, then, by the minimum degree condition, $e_{G^{\prime}}(X,Y)\geq\zeta n/12\cdot|X|\geq\zeta/12\cdot|X|\cdot|Y|$, as desired. So we may assume that $|X|,|Y|\geq\zeta n/12$. But then $e_{G^{\prime}}(X,Y)\geq\zeta/4\cdot|X|\cdot|Y|-4\varepsilon n^{2}\geq\zeta/12\cdot|X|\cdot|Y|$ since $G^{\prime}$ had no $\zeta/4$-sparse cuts (by (G’2)) and at most $4\varepsilon n^{2}$ edges are removed from $G^{\prime}$ by moving at most $4\varepsilon n$ vertices from $G^{\prime}$ to $V_{0}$. This shows that $G^{\prime}$ has no $\zeta/12$-sparse cuts, completing the proof of (G’8). Finally, to prove (G’9), note that if $G$ is bipartite and balanced, then by (G’4) and the above process of moving vertices, $G^{\prime}$ remains bipartite and balanced after moving the vertices. Otherwise, we claim that one needs to remove at least $\gamma n^{2}/4$ edges from $G^{\prime}$ to make it bipartite. Indeed, this follows from (G’3) and the fact that at most $4\varepsilon n^{2}\leq\gamma n^{2}/4$ edges are removed from $G^{\prime}$ after moving at most $4\varepsilon n$ vertices to $V_{0}$, proving (G’9).
5.2 Splitting the super-regular subgraphs
As shown in the previous subsection, the vertices outside $V_{0}$ can be covered by super-regular subgraphs $G^{\prime}[U_{2i-1},U_{2i}]$ for $i\in[m]$. In this subsection, we will show that if we randomly partition $U_{i}$ into five sets $\left\{U_{i}^{(1)},\ldots,U_{i}^{(5)}\right\}$ of appropriate sizes for each $i\in[2m]$, then the subgraph $H$ of $G^{\prime}$ induced by the set $\bigcup_{i\in[2m]}U_{i}^{(3)}$ still has good expansion properties (see (U6)) and, moreover, $H$ is either bipartite and balanced or is far from being bipartite (see (U7)).
Claim 5.5.
For every $i\in[2m]$, there is a partition $\left\{U_{i}^{(1)},\ldots,U_{i}^{(5)}\right\}$ of $U_{i}$ satisfying the following properties.
  1. (U1)
    $\big|U_{i}^{(1)}\big|=\big|U_{i}^{(2)}\big|=\xi|U_{i}|$ for $i\in[2m]$.
  2. (U2)
    $\big|U_{i}^{(3)}\big|=\frac{|U_{i}|}{2t}$ for $i\in[2m]$.
  3. (U3)
    $\big|U_{i}^{(4)}\big|=\frac{2|U_{i}|}{3}$ for $i\in[2m]$.
  4. (U4)
    $\big|U_{i}^{(5)}\big|=(\frac{1}{3}-\frac{1}{2t}-2\xi)|U_{i}|$ for $i\in[2m]$.
  5. (U5)
    For every $v\in V(G)$, $i\in[2m]$ and $j\in[5]$, if $d_{G}(v,U_{i})\geq\varepsilon|U_{i}|$, then $\frac{d_{G}\big(v,U_{i}^{(j)}\big)}{d_{G}(v,U_{i})}=(1\pm\varepsilon)\frac{|U_{i}^{(j)}|}{|U_{i}|}$.
    Similarly, for every $v\in V(G^{\prime})$, $i\in[2m]$ and $j\in[5]$, if $d_{G^{\prime}}(v,U_{i})\geq\varepsilon|U_{i}|$ then $\frac{d_{G^{\prime}}\big(v,U_{i}^{(j)}\big)}{d_{G^{\prime}}(v,U_{i})}=(1\pm\varepsilon)\frac{|U_{i}^{(j)}|}{|U_{i}|}$.
  6. (U6)
    The graph $H\coloneqq G^{\prime}\left[\bigcup_{i\in[2m]}U_{i}^{(3)}\right]$ has no $\frac{\zeta}{20}$-sparse cuts.
  7. (U7)
    If $G$ is bipartite and balanced, then $H=G^{\prime}\left[\bigcup_{i\in[2m]}U_{i}^{(3)}\right]$ is also bipartite and balanced. Otherwise, one needs to remove at least $\frac{\gamma}{256t^{2}}n^{2}$ edges from $H$ to make it bipartite.
For every $i\in[2m]$, let $\left\{U_{i}^{(1)},\ldots,U_{i}^{(5)}\right\}$ be a random partition of $U_{i}$ into sets of sizes $\xi|U_{i}|$, $\xi|U_{i}|$, $(1/2t)|U_{i}|$, $(2/3)|U_{i}|$, $(1/3-1/2t-2\xi)|U_{i}|$, respectively. Then (U1)(U4) hold by definition. Moreover, by Lemma 3.1 and a union bound, (U5) holds with probability at least 1o(1)1-o(1); fix an outcome such that it holds.
For proving (U6), let $\{X,Y\}$ be a partition of $V(H)$ with $|X|\leq|Y|$. Notice that by (U5) and (G’8), $H$ has minimum degree at least $(\zeta/7)\cdot|V(H)|$. Thus, if $|X|\leq(\zeta/14)\cdot|V(H)|$, then every vertex in $X$ has at least $(\zeta/14)\cdot|V(H)|$ neighbours in $Y$, showing that $e_{H}(X,Y)\geq|X|\cdot(\zeta/14)\cdot|V(H)|\geq(\zeta/14)\cdot|X|\cdot|Y|$, as required.
So suppose that $|X|\geq(\zeta/14)\cdot|V(H)|$. Let $\{X^{\prime},Y^{\prime}\}$ be a partition of $V(G^{\prime})$ such that, for every $i\in[2m]$, we have
$\frac{|X\cap U_{i}^{(3)}|}{|U_{i}^{(3)}|}=\frac{|X^{\prime}\cap U_{i}|}{|U_{i}|}.$ (15)
By (U2) and (15), it follows that $|X\cap U_{i}|=|X\cap U_{i}^{(3)}|=\frac{|X^{\prime}\cap U_{i}|}{2t}$ for every $i\in[2m]$. Thus, $|X|=\frac{|X^{\prime}|}{2t}$. Similarly, $|Y|=\frac{|Y^{\prime}|}{2t}$.
We claim that for any $i,j\in[2m]$, we have
$e_{H}(X\cap U_{i},Y\cap U_{j})\geq\frac{1}{4t^{2}}\cdot e_{G^{\prime}}(X^{\prime}\cap U_{i},Y^{\prime}\cap U_{j})-10\varepsilon|U_{i}||U_{j}|$ (16)
Indeed, if $|X^{\prime}\cap U_{i}|\leq 10t\varepsilon|U_{i}|$ or $|Y^{\prime}\cap U_{j}|\leq 10t\varepsilon|U_{j}|$, this holds trivially, as then the right-hand side of (16) is at most $(1/4t^{2})\cdot 10t\varepsilon|U_{i}||U_{j}|-10\varepsilon|U_{i}||U_{j}|<0$. If, instead, $|X^{\prime}\cap U_{i}|\geq 10t\varepsilon|U_{i}|$ and $|Y^{\prime}\cap U_{j}|\geq 10t\varepsilon|U_{j}|$, then $|X\cap U_{i}|=|X\cap U_{i}^{(3)}|\geq 5\varepsilon|U_{i}|$ and $|Y\cap U_{j}|=|Y\cap U_{j}^{(3)}|\geq 5\varepsilon|U_{j}|$ by (15) and (U2), so by $5\varepsilon$-regularity of $G^{\prime}[U_{i},U_{j}]$ (see (G’6)), we have
$\frac{e_{H}(X\cap U_{i},Y\cap U_{j})}{|X\cap U_{i}|\cdot|Y\cap U_{j}|}\geq\frac{e_{G^{\prime}}(X^{\prime}\cap U_{i},Y^{\prime}\cap U_{j})}{|X^{\prime}\cap U_{i}|\cdot|Y^{\prime}\cap U_{j}|}-10\varepsilon.$
This implies that
$e_{H}(X\cap U_{i},Y\cap U_{j})$ $\geq e_{G^{\prime}}(X^{\prime}\cap U_{i},Y^{\prime}\cap U_{j})\cdot\frac{|X\cap U_{i}|}{|X^{\prime}\cap U_{i}|}\cdot\frac{|Y\cap U_{j}|}{|Y^{\prime}\cap U_{j}|}-10\varepsilon\cdot|X\cap U_{i}|\cdot|Y\cap U_{j}|$
[…]
as required for (16).
Hence, we obtain the following lower bound on the number of edges between $X$ and $Y$ in $H$.
$e_{H}(X,Y)$ $=\sum_{i,j\in[2m]}e_{H}(X\cap U_{i},Y\cap U_{j})$
[…]
$=\frac{1}{4t^{2}}\cdot e_{G^{\prime}}(X^{\prime},Y^{\prime})-10\varepsilon\cdot|V(G^{\prime})|^{2}$
[…]
$=\frac{\zeta}{12}\cdot|X|\cdot|Y|-10\varepsilon n^{2}$
$\geq\frac{\zeta}{20}\cdot|X|\cdot|Y|,$
using that $|X|,|Y|\geq(\zeta/14)|V(H)|\geq(\zeta/56t)n$ and that $\varepsilon\ll 1/t,\zeta$ for the last inequality. This proves (U6).
Finally, it remains to prove (U7). Note that by (G’4) and (U2), if $G$ is bipartite and balanced, then $H$ is also bipartite and balanced, as desired. Now suppose that $G$ is $\gamma$-far-from-bipartite. Then, by (G’9), one needs to remove at least $\gamma n^{2}/4$ edges from $G^{\prime}$ to make it bipartite. Let $\{X,Y\}$ be an arbitrary partition of $V(H)$. To prove (U7), we need to show that $e_{H}(X)+e_{H}(Y)\geq\frac{\gamma n^{2}}{256t^{2}}$. Define the sets $X^{\prime},Y^{\prime}$ as in (15). Then, analogously to (16), for every $i,j\in[2m]$, we have
$e_{H}(X\cap U_{i},X\cap U_{j})\geq\frac{1}{4t^{2}}\cdot e_{G^{\prime}}(X^{\prime}\cap U_{i},X^{\prime}\cap U_{j})-10\varepsilon|U_{i}||U_{j}|.$ (17)
It follows that
$e_{H}(X)$ $=\sum_{1\leq i<j\leq 2m}e_{H}(X\cap U_{i},X\cap U_{j})$
[…]
$\geq\frac{1}{4t^{2}}\cdot e_{G^{\prime}}(X^{\prime})-10\varepsilon n^{2}.$
Analogously, $e_{H}(Y)\geq\frac{1}{4t^{2}}\cdot e_{G^{\prime}}(Y^{\prime})-10\varepsilon n^{2}$. Notice that $e_{G^{\prime}}(X)+e_{G^{\prime}}(Y)\geq\gamma n^{2}/4$ by (G’9) and the assumption that $G$ is not bipartite. Hence, altogether, we have
$e_{H}(X)+e_{H}(Y)\geq\frac{1}{4t^{2}}\big(e_{G^{\prime}}(X^{\prime})+e_{G^{\prime}}(Y^{\prime})\big)-20\varepsilon n^{2}\geq\frac{\gamma}{16t^{2}}\cdot n^{2}-20\varepsilon\cdot n^{2}\geq\frac{\gamma}{256t^{2}}\cdot n^{2},$
using $\varepsilon\ll\gamma,1/t$ for the last inequality. This proves (U7). ∎
5.3 Covering $V_{0}$ using a well-distributed collection of copies of $K_{t,t}$
In the rest of the proof, let $U^{(j)}\coloneqq\bigcup_{i\in[2m]}U_{i}^{(j)}$ for $j\in[5]$. In this subsection, we find a collection of vertex-disjoint copies of $K_{t,t}$ in $G\left[V_{0}\cup U^{(1)}\cup U^{(2)}\right]$ that covers $V_{0}$ such that the number of uncovered vertices in each $U_{i}$ is divisible by $2t$. In the rest of the proof, using a slight abuse of notation, we denote the set of vertices contained in a given collection $\mathcal{K}$ of copies of $K_{t,t}$ in $G$ by $V(\mathcal{K})$.
First, we cover the vertices in $V_{0}$ using vertex-disjoint $K_{t,t}$’s in $G\left[V_{0}\cup U^{(1)}\right]$ (ignoring the divisibility requirement).
Claim 5.6.
There is a collection $\mathcal{K}_{1}$ of pairwise vertex-disjoint $K_{t,t}$’s in $G[V_{0}\cup U^{(1)}]$ such that $V_{0}\subseteq V(\mathcal{K}_{1})$ and each copy of $K_{t,t}$ in $\mathcal{K}_{1}$ contains exactly one vertex of $V_{0}$.
Let $\mathcal{K}_{1}$ be a maximal collection of pairwise vertex-disjoint copies of $K_{t,t}$ in $G[V_{0}\cup U^{(1)}]$, where each copy of $K_{t,t}$ contains exactly one vertex from $V_{0}$. We will show that $V_{0}\subseteq V(\mathcal{K}_{1})$, proving the claim. Suppose not, and let $v\in V_{0}-V(\mathcal{K}_{1})$. Write $U\coloneqq U^{(1)}$ and $U^{\prime}\coloneqq U-V(\mathcal{K}_{1})$.
We claim that every vertex $u$ in $G$ satisfies $d_{G}(u,U^{\prime})\geq(\xi\zeta/8)n$. Indeed, given a vertex $u$ in $G$, let $I$ be the set of indices $i\in[2m]$ such that $d_{G}(u,U_{i})\geq\varepsilon|U_{i}|$. Then
[…]
$\geq(1-\varepsilon)\cdot\xi\cdot\left(d_{G}(u)-d_{G}(u,V_{0})-\sum_{i\in[2m]-I}d_{G}(u,U_{i})\right)$
$\geq(1-\varepsilon)\cdot\xi\cdot\left(\frac{\zeta n}{2}-|V_{0}|-\sum_{i\in[2m]-I}\varepsilon|U_{i}|\right)$
[…]
For the third inequality we used that $G$ has minimum degree at least $\zeta(n-1)\geq\frac{\zeta}{2}n$, which follows from the fact that there are no $\zeta$-sparse cuts in $G$. Since $|V(\mathcal{K}_{1})|\leq 2t|V_{0}|\leq 12\varepsilon tn$, every vertex $u$ in $G$ satisfies $d_{G}(u,U^{\prime})\geq d_{G}(u,U)-|V(\mathcal{K}_{1})|\geq(\frac{\xi\zeta}{4}-12\varepsilon t)n\geq(\xi\zeta/8)n$, as desired.
Fix a subset $W\subseteq N(v,U^{\prime})$ of size $(\xi\zeta/16)n$ and write $Z\coloneqq U^{\prime}-W$. Since every vertex $u$ in $G$ satisfies $d_{G}(u,U^{\prime})\geq(\xi\zeta/8)n$, every vertex in $W$ has at least $(\xi\zeta/16)n$ neighbours in $Z$. By Lemma 3.8, the graph $G[W,Z]$ has a $K_{t,t}$, say, with vertex set $A\cup B$ where $A\subseteq W$ and $B\subseteq Z$. Let $B^{\prime}$ be a subset of $B$ of size t1t-1. Then $G[A,B^{\prime}\cup\{v\}]$ is a copy of $K_{t,t}$ in $G[V_{0}\cup U^{(1)}]$, which is vertex-disjoint from $V(\mathcal{K}_{1})$ and contains exactly one vertex from $V_{0}$, contradicting the maximality of $\mathcal{K}_{1}$. ∎
5.4 Constructing copies of $K_{t,t}$ to ensure divisibility
Next, we find a small number of vertex-disjoint $K_{t,t}$’s in $G[U^{(2)}]$ that will allow us to ensure that the number of uncovered vertices in every $U_{i}$ is divisible by $t$. Recall that the support of a function $f:A\to B$, denoted $\operatorname{supp}(f)$, is the set of elements $a$ in $A$ with $f(a)\neq 0$.
Claim 5.7.
Let $f:[2m]\to\{0,\ldots,t-1\}$ be a function such that $\sum_{i\in[2m]}f(i)$ is divisible by $t$, and if $G$ is bipartite, then $\sum_{i\in[m]}f(2i)$ is also divisible by $t$. Then there is a collection $\mathcal{K}_{2}$ of pairwise vertex-disjoint $K_{t,t}$’s in $G[U^{(2)}]$ such that $|U_{i}\cap V(\mathcal{K}_{2})|\leq 3t\cdot|\operatorname{supp}(f)|$ and $|U_{i}\cap V(\mathcal{K}_{2})|\equiv f(i)\pmod{t}$ for $i\in[2m]$.
We prove the claim by induction on the size of $\operatorname{supp}(f)$. If $|\operatorname{supp}(f)|=0$, then we can take $\mathcal{K}_{2}=\emptyset$. So, suppose that $\operatorname{supp}(f)\neq\emptyset$, and let $a\in[2m]$ satisfy $f(a)\neq 0$. By the assumption that $\sum_{i\in[2m]}f(i)$ is divisible by $t$, there is $b\in[2m]-\{a\}$ such that $f(b)\neq 0$, and, if $G$ is bipartite, we may take $b$ to have the same parity as $a$.
Define $f^{\prime}:[2m]\to\{0,\ldots,t-1\}$ as follows.
$f^{\prime}(i)=\left\{\begin{array}[]{ll}f(i)&i\neq a,b,\\ 0&i=a,\\ f(a)+f(b)&i=b.\end{array}\right$.
Notice that $\sum_{i\in[2m]}f^{\prime}(i)$ is divisible by $t$, $\sum_{i\in[m]}f^{\prime}(2i)$ is also divisible by $t$ if $G$ is bipartite and $|\operatorname{supp}(f^{\prime})|<|\operatorname{supp}(f)|$. Thus, by the induction hypothesis, there is a collection $\mathcal{K}^{\prime}$ of pairwise vertex-disjoint $K_{t,t}$’s in $G[U^{(2)}]$ such that $|U_{i}\cap V(\mathcal{K}^{\prime})|\leq 3t\cdot|\operatorname{supp}(f^{\prime})|$ and $|U_{i}\cap V(\mathcal{K}^{\prime})|\equiv f^{\prime}(i)\!\pmod{t}$ for $i\in[2m]$.
We claim that there is a walk $W$ of even length from $U_{a}$ to $U_{b}$ in $\Gamma_{\mathrm{ref}}$ that does not visit a vertex more than three times. Indeed, notice that $\Gamma_{\mathrm{ref}}$ is connected since $G^{\prime}$ has no $\zeta/12$-sparse cuts (see (G’8)). If $G$ is bipartite, any path in $\Gamma_{\mathrm{ref}}$ from $U_{a}$ to $U_{b}$ suffices: such a path exists by connectivity, and because $a$ and $b$ have the same parity, its length is necessarily even. If $G$ is not bipartite, then (G’9) (and the fact that $\mu\ll\gamma$) implies that $\Gamma_{\mathrm{ref}}$ is also not bipartite. Let $C$ be an odd cycle in $\Gamma_{\mathrm{ref}}$, and let $v\in V(C)$. By connectivity of $\Gamma_{\mathrm{ref}}$, there are paths $P_{a}$ and $P_{b}$ from $a$ to $v$ and from $b$ to $v$. If the lengths of $P_{a}$ and $P_{b}$ have the same parity, then $aP_{a}vP_{b}b$ is a walk of even length visiting each vertex at most twice. Otherwise, $aP_{a}vCvP_{b}b$ is a walk of even length that visits each vertex at most three times, as desired.
Let $W$ be a walk in $\Gamma_{\mathrm{ref}}$ as in the previous paragraph, and write $W=U_{i_{0}},\ldots,U_{i_{2\ell}}$, where $U_{i_{0}}=U_{a}$ and $U_{i_{2\ell}}=U_{b}$. Since $W$ visits each vertex in $\Gamma_{\mathrm{ref}}$ at most three times and $|U_{i}\cap V(\mathcal{K}^{\prime})|\leq 3t\cdot|\operatorname{supp}(f^{\prime})|\leq 6tm$ for $i\in[2m]$, we can apply Lemma 3.8 to iteratively choose pairwise vertex-disjoint $K_{t,t}$’s in $U^{(2)}-V(\mathcal{K}^{\prime})$, denoted $K_{1},\ldots,K_{\ell}$, such that for every $j\in[\ell]$, $K_{j}$ consists of $t$ vertices in $U^{(2)}_{i_{2j-1}}$, $f(a)$ vertices in $U^{(2)}_{i_{2j-2}}$ and tf(a)t-f(a) vertices in $U^{(2)}_{i_{2j}}$. Write $\mathcal{K}_{2}\coloneqq\mathcal{K}^{\prime}\cup\{K_{1},\ldots,K_{\ell}\}$. It is easy to check that $\mathcal{K}_{2}$ satisfies the requirements of the induction hypothesis since $|U_{i}\cap V(\mathcal{K}_{2})|\leq 3t\cdot|\operatorname{supp}(f^{\prime})|+3t\leq 3t|\operatorname{supp}(f)|$ and $|U_{i}\cap V(\mathcal{K}_{2})|\equiv f(i)\!\pmod{t}$ for $i\in[2m]$. This completes the induction step and proves the claim. ∎
Let $\mathcal{K}_{1}$ be a collection of copies of $K_{t,t}$ as guaranteed by Claim 5.6. For $i\in[2m]$, let $f(i)\coloneqq-|U_{i}\cap V(\mathcal{K}_{1})|\pmod{t}=|U_{i}^{(1)}\cap V(\mathcal{K}_{1})|\pmod{t}$. Then
$\sum_{i\in[2m]}f(i)\equiv\sum_{i\in[2m]}|U_{i}\cap V(\mathcal{K}_{1})|\equiv|V(\mathcal{K}_{1})|-|V_{0}|\equiv 0\pmod{2t}$, (18)
where the second equality uses that $V(G)=V_{0}\cup U_{1}\cup\cdots\cup U_{2m}$ and $V_{0}\subseteq V(\mathcal{K}_{1})$, and the third follows from the fact that both $|V(\mathcal{K}_{1})|$ and $|V_{0}|$ are divisible by $2t$, with the divisibility of $|V_{0}|$ given by (G’4).
If $G$ is bipartite (and balanced), then, denoting by $\{X,Y\}$ the bipartition of $G$ with $X-V_{0}=\bigcup_{i\in[m]}U_{2i-1}$ and $Y-V_{0}=\bigcup_{i\in[m]}U_{2i}$,
$\sum_{i\in[m]}f(2i)\equiv\sum_{i\in[m]}|U_{2i}\cap V(\mathcal{K}_{1})|\equiv|(Y-V_{0})\cap V(\mathcal{K}_{1})|\equiv\frac{|V(\mathcal{K}_{1})|-|V_{0}|}{2}\equiv 0\pmod{t}$. (19)
The first equality follows by the definition of $f$, and the second equality follows from the fact that $Y-V_{0}=\bigcup_{i\in[m]}U_{2i}$. For the third equality, first recall that $|U_{1}|=\ldots=|U_{2m}|$ by (G’4), so $|X-V_{0}|=|Y-V_{0}|$. This implies that $|X\cap V_{0}|=|Y\cap V_{0}|$ since $G$ is bipartite and balanced. Combining this with the fact that $V_{0}\subseteq V(\mathcal{K}_{1})$, and that every copy of $K_{t,t}$ in $\mathcal{K}_{1}$ contains exactly one vertex of $V_{0}$, we get $|(X-V_{0})\cap V(\mathcal{K}_{1})|=|(Y-V_{0})\cap V(\mathcal{K}_{1})|$ from which the third equality follows easily. The final equality follows from $|V_{0}|$ and $|V(\mathcal{K}_{1})|$ being divisible by $2t$. Using (18), (19) and Claim 5.7, we can find a collection $\mathcal{K}_{2}$ of pairwise vertex-disjoint copies of $K_{t,t}$ in $G[U^{(2)}]$ satisfying $|U_{i}^{(2)}\cap V(\mathcal{K}_{2})|\equiv f(i)\equiv-|U_{i}^{(1)}\cap V(\mathcal{K}_{1})|\pmod{t}$ for $i\in[2m]$, which implies that $|U_{i}\cap V(\mathcal{K}_{1}\cup\mathcal{K}_{2})|$ is divisible by $t$, as desired.
5.5 Covering the remaining vertices of $G$ with vertex-disjoint copies of $K_{t,t}$
In this subsection, our aim is to cover the remaining vertices of $G$ with vertex-disjoint copies of $K_{t,t}$. To this end, we will find a collection $\mathcal{K}_{3}$ of copies of $K_{t,t}$ such that $V(\mathcal{K}_{3})\subseteq V(G^{\prime})-V(\mathcal{K}_{1}\cup\mathcal{K}_{2})$ and such that the sets $U^{\prime}_{i}\coloneqq U_{i}-V(\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3})$, $i\in[2m]$, have the same size. This allows us to apply the blow-up lemma to find perfect $K_{t,t}$-packings in $G^{\prime}[U^{\prime}_{2i-1},U^{\prime}_{2i}]$ for every $i\in[m]$, thereby covering all of the remaining vertices, as desired.
Let us now carry out the above strategy. For $i\in[2m]$, let $U_{i}^{\prime(3)}$ be an arbitrary subset of $U_{i}^{(3)}$ such that
$\big|U_{i}^{\prime(3)}\big|=\big|U_{i}^{(3)}\big|-\frac{\big|U_{i}\cap V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\big|}{t}.$ (20)
(Notice that the right-hand-side of (20) is an integer due to the choice of $\mathcal{K}_{2}$.) Recall that $U^{(3)}=\bigcup_{i\in[2m]}U_{i}^{(3)}$. Write $U^{\prime(3)}\coloneqq\bigcup_{i\in[2m]}U^{\prime(3)}_{i}$. In the rest of this section, let $H\coloneqq G^{\prime}\left[U^{(3)}\right]$, $H^{\prime}\coloneqq G^{\prime}\left[U^{\prime(3)}\right]$ and let $d^{\prime}\coloneqq d/2t$.
Claim 5.8.
There is a perfect matching in $H^{\prime}$.
We claim that the following properties hold.
  1. (a)
    $H$ has maximum degree at most $d^{\prime}(1+\eta)$.
  2. (b)
    $H$ has average degree at least $d^{\prime}(1+\eta)-32\eta t\cdot|V(H)|$.
  3. (c)
    $H$ has no $\zeta/20$-sparse cuts.
  4. (d)
    $H$ is either bipartite and balanced, or one needs to remove at least $\frac{\gamma}{256t^{2}}n^{2}$ edges from $H$ to make it bipartite.
  5. (e)
    $|V(H)-V(H^{\prime})|\leq 8t\xi\cdot|V(H)|$.
  6. (f)
    If $G$ is bipartite and balanced, then $H^{\prime}$ is also bipartite and balanced.
Indeed, by (U2) and (U5), any vertex $v$ in $H$ has degree at most $\sum_{i\in[2m]}\left((1+\varepsilon)\frac{d_{G^{\prime}}(v,U_{i})}{2t}\right)+\varepsilon n\leq d^{\prime}(1+\varepsilon)+\varepsilon n\leq d^{\prime}(1+\eta)$, proving (a). For proving (b), we have the following sequence of inequalities.
[…]
$=\frac{1-\varepsilon}{2t}\left(\sum_{v\in V(G^{\prime})}d_{G^{\prime}}(v)-\sum_{v\in V(G^{\prime})-V(H)}d_{G^{\prime}}(v)\right)$
[…]
$=\frac{1-\varepsilon}{2t}\cdot\Big(d\cdot|V(H)|-6\eta n\cdot|V(G^{\prime})|\Big)$
$\geq d^{\prime}\cdot|V(H)|-\varepsilon d^{\prime}\cdot|V(H)|-6\eta n\cdot|V(H)|$
$=d^{\prime}(1+\eta)\cdot|V(H)|-(\varepsilon d^{\prime}+6\eta n+\eta d^{\prime})\cdot|V(H)|$
$\geq d^{\prime}(1+\eta)\cdot|V(H)|-8\eta n\cdot|V(H)|$
$\geq d^{\prime}(1+\eta)\cdot|V(H)|-32\eta t\cdot|V(H)|^{2}.$
It follows that $H$ has average degree at least $d^{\prime}(1+\eta)-32\eta t\cdot|V(H)|$, as required for (b). Notice that (c) is the same as (U6) and (d) is the same as (U7). For proving (e), note that $|V(H)-V(H^{\prime})|\leq|U^{(1)}|+|U^{(2)}|\leq 2\xi n\leq 8t\xi\cdot|V(H)|$. Finally, for proving (f), suppose that $G$ is bipartite and balanced, and denote its bipartition by $\{X,Y\}$. Recall that, by (G’4), $|V_{0}\cap X|=|V_{0}\cap Y|$ (since $|X|=|Y|$). By the choice of $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ (see Claim 5.6 and Claim 5.7), we have $|V(G^{\prime})\cap(V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap X)|=|V(G^{\prime})\cap(V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap Y)|$, and thus
$|V(H^{\prime})\cap X|$ $=\sum_{i\in[m]}\left(|U_{2i-1}^{(3)}|-\frac{|U_{2i-1}\cap V(\mathcal{K}_{1}\cup\mathcal{K}_{2})|}{t}\right)$
[…]
$=\sum_{i\in[m]}\left(|U_{2i}^{(3)}|-\frac{|U_{2i}\cap V(\mathcal{K}_{1}\cup\mathcal{K}_{2}|)}{t}\right)=|V(H^{\prime})\cap Y|,$
proving that $H^{\prime}$ is balanced, as required for (f).
Thus, by applying Lemma 3.3 with $H$, $d^{\prime}(1+\eta)$, $c/2t$, $32\eta t$, $8t\xi$, $\gamma/(256t^{2})$, $\zeta/20$, $V(H^{\prime}), V(H)V(H)$ playing the roles of $G$, $d$, $c$, $\eta$, $\xi$, $\gamma$, $\zeta$, $W$, $n$ respectively, we find that there is a Hamilton cycle in $H^{\prime}$, which contains a perfect matching in $H^{\prime}$ (since $|V(H^{\prime})|$ is even), thus proving the claim. ∎ Let $\mathcal{M}^{\prime}$ be a perfect matching in $H^{\prime}$ (guaranteed by Claim 5.8). For $i,j\in[2m]$, denote by $f(i,j)$ the number of edges in $\mathcal{M}^{\prime}[U_{i}^{\prime(3)},U_{j}^{\prime(3)}]$.
Claim 5.9.
There is a collection $\mathcal{K}_{3}$ of pairwise vertex-disjoint copies of $K_{t,t}$ in $G^{\prime}\left[U^{(4)}\right]$ such that for every $i,j\in[2m]$, $\mathcal{K}_{3}$ consists of $f(i,j)$ copies of $K_{t,t}$ in $G^{\prime}[U_{i},U_{j}]$ (and there are no other $K_{t,t}$’s).
Let $\mathcal{K}_{3}$ be a maximal collection of pairwise vertex-disjoint $K_{t,t}$’s in $G^{\prime}\left[U^{(4)}\right]$, such that for every $i,j\in[2m]$, $\mathcal{K}_{3}$ consists of at most $f(i,j)$ copies of $K_{t,t}$ in $G^{\prime}[U_{i}^{(4)},U_{j}^{(4)}]$ (and no other copies of $K_{t,t}$). Suppose towards a contradiction that there exist $i,j\in[2m]$ such that the number of copies of $K_{t,t}$ in $\mathcal{K}_{3}$ that are in $G^{\prime}[U_{i},U_{j}]$ is strictly smaller than $f(i,j)$ (implying that $f(i,j)>0$). Write $U_{\ell}^{\prime\prime}\coloneqq U_{\ell}^{(4)}-V(\mathcal{K}_{3})$ for every $\ell\in[2m]$.
We claim that $|U_{\ell}^{\prime\prime}|\geq|U_{\ell}|/6$ for every $\ell\in[2m]$. Indeed, note that for every $\ell\in[2m]$ we have $|U_{\ell}^{(4)}\cap V(\mathcal{K}_{3})|\leq t\cdot\sum_{j\in[2m]}f(\ell,j)=t\cdot|U_{\ell}^{\prime(3)}|\leq t\cdot|U_{\ell}^{(3)}|=|U_{\ell}|/2$, where in the last equality we used (U2). Thus, by (U3), for every $\ell\in[2m]$, $|U_{\ell}^{\prime\prime}|=|U_{\ell}^{(4)}|-|U_{\ell}^{(4)}\cap V(\mathcal{K}_{3})|\geq(2/3)|U_{\ell}|-(1/2)|U_{\ell}|=|U_{\ell}|/6$, as claimed.
Now recall that, by (G’6), for all $i,j\in[2m]$, the graph $G^{\prime}[U_{i},U_{j}]$ is $5\varepsilon$-regular, and it has density more than $\mu/5$ if it has at least one edge. Hence, we have $e_{G^{\prime}}(U_{i}^{\prime\prime},U_{j}^{\prime\prime})>(\mu/5-5\varepsilon)\cdot|U_{i}^{\prime\prime}|\cdot|U_{j}^{\prime\prime}|$, so by the Kővári–Sós–Turán theorem (Lemma 3.8), there is a $K_{t,t}$ in $G^{\prime}[U_{i}^{\prime\prime},U_{j}^{\prime\prime}]$, contradicting the maximality of $\mathcal{K}_{3}$ and proving the claim. ∎
Let $\mathcal{K}_{3}$ be a collection of copies of $K_{t,t}$ as guaranteed by Claim 5.9. For $i\in[2m]$, write $U_{i}^{\prime}\coloneqq U_{i}-V(\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3})$. For $i\in[2m]$, by the choice of $U_{i}^{\prime(3)}$ as in (20) and the fact that $|U_{i}\cap V(\mathcal{K}_{3})|=t\cdot\sum_{j\in[2m]}f(i,j)=t\cdot|U_{i}^{\prime(3)}|$, we have
$|U_{i}^{\prime}|$ $=|U_{i}-V(\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3})|=|U_{i}|-\big|U_{i}\cap V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\big|-\big|U_{i}\cap V(\mathcal{K}_{3})\big|$
$=|U_{i}|-\big|U_{i}\cap V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\big|-t\cdot\left(\big|U_{i}^{(3)}\big|-\frac{\big|U_{i}\cap V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\big|}{t}\right)=|U_{i}|-t\big|U_{i}^{(3)}\big|=\frac{|U_{i}|}{2}.$
Recall that the sets $U_{i}$ for $i\in[2m]$ have the same size, which is divisible by $2t$ by (G’4). Hence, the sets $U^{\prime}_{i}$ for $i\in[2m]$ have the same size, which is divisible by $t$. Moreover, by (U4), (U5), (G’5) and the fact that $U_{i}^{(5)}$ is a subset of $U_{i}^{\prime}$ for each $i\in[2m]$, we have that $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ is $(10\varepsilon,\mu/50)$-super-regular for $i\in[m]$. Therefore, it follows from the blow-up lemma (Lemma 3.6) that $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ has a perfect $K_{t,t}$-packing (since $\varepsilon\ll\mu,1/t$). These perfect $K_{t,t}$-packings of $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ for $i\in[m]$, along with $\mathcal{K}_{1}$, $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$, form a perfect $K_{t,t}$-packing in $G$. This completes the proof of Lemma 5.1. ∎
6 $F$-packings in dense regular graphs
In this section, we combine Lemma 4.1 and Corollary 5.2 to prove Theorem 1.4, which asserts that for every bipartite graph $F$ and every constant $0<c\leq 1$, there is a constant $C>0$ such that every $d$-regular graph $G$ of order $n$, with $d\geq cn$, has an $F$-packing that covers all but at most $C$ vertices of $G$. Denote the sizes of the two parts of $F$ by $a$ and $b$, and write $t\coloneqq a+b$. Then, it is easy to see that there are two vertex-disjoint copies of $F$ covering all the vertices of the complete bipartite graph $K_{t,t}$. Hence, Theorem 1.4 follows immediately from the following.
Theorem 6.1.
Let $0<c\leq 1$, and let $t\geq 2$ be an integer. Then there exists a constant $C=C(t,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has a $K_{t,t}$-packing that covers all but at most $C$ vertices of $G$.
Let $n_{0}\in\mathbb{N}$ be chosen sufficiently large depending on $c$ and $t$, that is, $1/n_{0}\ll c$. If $n<n_{0}$, then the conclusion holds trivially for a suitably large constant $C=C(t,c)$. So, assume $n\geq n_{0}$. Then by applying Lemma 4.1 to $G$, we obtain positive numbers $\eta,\gamma,\zeta,r$ with $r\leq\lceil 1/c\rceil$, a subgraph $G^{\prime}\subseteq G$ with $V(G^{\prime})=V(G)$ and pairwise disjoint sets $Z_{1},\ldots,Z_{r},L\subseteq V(G)$ satisfying properties (B1)(B5) such that
$1/n_{0}\ll\eta\ll\zeta\ll\gamma\ll c,1/t.$
For all $i\in[r]$, by (B2), $G^{\prime}[Z_{i}]$ has average degree at least $d-\eta|Z_{i}|$, by (B3), $G^{\prime}[Z_{i}]$ has no $\zeta$-sparse cuts and, by (B4), either $G^{\prime}[Z_{i}]$ is bipartite and balanced, or one needs to remove at least $\gamma n^{2}\geq\gamma|Z_{i}|^{2}$ edges to make it bipartite, and $d\geq cn\geq c|Z_{i}|$. Thus, for every $i\in[r]$, we can apply Corollary 5.2 with $G^{\prime}[Z_{i}]$ playing the role of $G$, to obtain a $K_{t,t}$-packing of $G^{\prime}[Z_{i}]$ covering all but at most 2t12t-1 vertices in $Z_{i}$. Moreover, $G[L]$ has a perfect $K_{t,t}$-packing by (B5), and $|V(G)-(Z_{1}\cup\ldots\cup Z_{r}\cup L)|\leq 64r^{2}\cdot\frac{t}{c}\cdot\left(\frac{e}{8\zeta}\right)^{t}$ by (B1). This yields a $K_{t,t}$-packing of $G$ covering all but at most the following number of vertices.
$r\cdot(2t-1)+64\cdot r^{2}\cdot\frac{t}{c}\cdot\left(\frac{e}{8\zeta}\right)^{t}\leq 66\cdot r^{2}\cdot\frac{t}{c}\cdot\left(\frac{e}{8\zeta}\right)^{t}\leq 66\left(\left\lceil\frac{1}{c}\right\rceil\right)^{2}\cdot\frac{t}{c}\cdot\left(\frac{e}{8\zeta}\right)^{t}.$
This completes the proof of Theorem 6.1. ∎
7 Packing subdivisions in dense regular graphs
In this section, we prove Theorem 1.7. We start by giving a detailed sketch of its proof here. Let $F$ be a graph, and let $0\leq c\leq 1$. Let $G$ be a $d$-regular graph of order $n$, where $d\geq cn$. We start, as before, by partitioning the vertex set of $G$ into sets $Z_{1},\ldots,Z_{r}$ such that the subgraphs $G[Z_{i}]$, $i\in[r]$ induced by these sets are expanders with few edges between them (using Lemma 3.2). Let $\{X_{i},Y_{i}\}$ be a bipartition of $G$ maximising the number of edges in $G[X_{i},Y_{i}]$. Let $G^{\prime}$ be the graph obtained from $G$ by removing all edges within $X_{i}$ and $Y_{i}$ for all $i\in[r]$ such that $G[Z_{i}]$ is close to being bipartite. This ensures that for all $i\in[r]$, $G^{\prime}[Z_{i}]$ is either bipartite with the bipartition $\{X_{i},Y_{i}\}$ (with roughly the same average degree and minimum degree as $G[Z_{i}]$) or it is far from being bipartite.
Next, we find short paths $P_{1},\ldots,P_{r}$ (where each $P_{i}$ has both of its ends in $Z_{i}$) whose removal ‘balances’ the expanders that are close to being bipartite. More precisely, for every $i\in[r]$, if $G[Z_{i}]$ is close to being bipartite, then, writing $Q\coloneqq P_{1}\cup\ldots\cup P_{r}$, we have $|X_{i}-V(Q)|=|Y_{i}-V(Q)|$, and $P_{i}$ has one end in $X_{i}$ and the other end in $Y_{i}$. The key tool for building these paths is Lemma 3.4 due to Gruslys and Letzter [11], which produces a small linear forest $H$ that balances the expanders that are close to being bipartite, while also ensuring that each expander contains zero or two leaves of $H$. To obtain the desired paths $P_{1},\ldots,P_{r}$, we iteratively merge pairs of components of $H$ whose leaves lie in the same expander and, if necessary, we define some paths $P_{i}$ as arbitrary edges within $G^{\prime}[Z_{i}]$.
Our next step is to find two small, vertex-disjoint $F$-subdivisions within each expander $G^{\prime}[Z_{i}]$, denoted $F_{i}$ and $F_{i}^{\prime}$, which are disjoint from $Q$ and whose union is balanced; that is, if $G[Z_{i}]$ is close to being bipartite, then $|X_{i}\cap V(F_{i}\cup F_{i}^{\prime})|=|Y_{i}\cap V(F_{i}\cup F_{i}^{\prime})|$. We construct $F_{i}$ and $F_{i}^{\prime}$ greedily: we begin by selecting $|V(F)|$ vertices from $X_{i}$ to serve as the branch vertices of $F_{i}$, and $|V(F)|$ vertices from $Y_{i}$ to serve as the branch vertices of $F_{i}^{\prime}$. We then iteratively connect pairs of branch vertices (corresponding to the edges of $F$) one pair at a time, to complete each subdivision.
Finally, for each $i\in[r]$, we absorb the path $P_{i}$ and the remaining uncovered vertices in $Z_{i}$ into the subdivision $F_{i}$. To do this, we replace an arbitrary edge $x_{i}y_{i}$ in $F_{i}$ with a path that starts at $x_{i}$, connects to one end of $P_{i}$, traverses $P_{i}$, and then continues through all the remaining uncovered vertices in $Z_{i}$ before returning to $y_{i}$.
In the last three steps, we rely on the fact that the expanders $G^{\prime}[Z_{i}]$ are robustly connected via short paths; that is, any two vertices can be joined by a short path that avoids any given small set of forbidden vertices (see Lemma 7.1). In the final step, we also use the robust Hamiltonicity property of the expanders $G^{\prime}[Z_{i}]$, meaning that for any small set of forbidden vertices $W$ (which is balancing if $G^{\prime}[Z_{i}]$ is bipartite), and any two vertices $x,y\in Z_{i}-W$ (lying in different parts of $G^{\prime}[Z_{i}]$ if it is bipartite), there exists a Hamilton path in $G^{\prime}[Z_{i}]-W$ with ends $x$ and $y$.
The proof of Theorem 1.7 is given in Section 7.3, after establishing two preparatory lemmas in the following two subsections.
7.1 Expanders are robustly connected via short paths
In the following lemma, we show that our expanders are ‘robustly connected’ via short paths, a property we will use several times in the proof.
Lemma 7.1.
Let $n$ be a positive integer, and let $\beta,\xi,\zeta,\delta\in(0,1)$ satisfy $1/n\ll\xi,\beta\ll\zeta\ll\delta$. Let $H$ be a graph on at most $n$ vertices with minimum degree at least $\delta n$ and no $\zeta$-sparse cuts. If $H$ is $\beta$-almost-bipartite, let $H^{\prime}$ be a largest bipartite subgraph of $H$; otherwise, set $H^{\prime}:=H$. Then, $H^{\prime}$ has no $\zeta/2$-sparse cuts. Moreover, if $W\subseteq V(H)$ is a subset of size at most $\xi n$, then $H^{\prime\prime}\coloneqq H^{\prime}-W$ satisfies the following properties.
  1. (i)
    $H^{\prime\prime}$ has no $\zeta/2$-sparse cuts.
  2. (ii)
    For every pair $p,p^{\prime}\in V(H^{\prime\prime})$, there exists a $(p,p^{\prime})$-path in $H^{\prime\prime}$ of length at most $15/\delta$.
Notice that $H^{\prime}$ has minimum degree at least $\delta n/2$, and $H^{\prime\prime}$ has minimum degree at least $\delta n/4$. First, we show that $H^{\prime}$ has no $\zeta/2$-sparse cuts. Consider a partition $\{X,Y\}$ of $V(H^{\prime})$ with $|X|\leq|Y|$. If $|X|\leq\delta n/4$, then, since $H^{\prime}$ has minimum degree at least $\delta n/2$, we have $e_{H^{\prime}}(X,Y)\geq\frac{\delta n}{4}\cdot|X|\geq\frac{\delta}{4}\cdot|X|\cdot|Y|\geq\frac{\zeta}{2}\cdot|X|\cdot|Y|$. So we may assume that $|X|,|Y|\geq\delta n/4$. Then, since $H$ has no $\zeta$-sparse-cuts, we have $e_{H^{\prime}}(X,Y)\geq e_{H}(X,Y)-\beta n^{2}\geq\zeta\cdot|X|\cdot|Y|-\beta n^{2}\geq\zeta/2\cdot|X|\cdot|Y|$, showing that $H^{\prime}$ has no $\zeta/2$-sparse cuts, as desired. To prove (i), we use a very similar argument. Consider a partition $\{X,Y\}$ of $V(H^{\prime\prime})$ with $|X|\leq|Y|$. If $|X|\leq\delta n/8$, then, since $H^{\prime\prime}$ has minimum degree at least $\delta n/4$, we have $e_{H^{\prime\prime}}(X,Y)\geq\frac{\delta n}{8}\cdot|X|\geq\frac{\delta}{8}\cdot|X|\cdot|Y|\geq\frac{\zeta}{2}\cdot|X|\cdot|Y|$. So we may assume that $|X|,|Y|\geq\delta n/8$. But then, since $H$ has no $\zeta$-sparse-cuts,
$e_{H^{\prime\prime}}(X,Y)=e_{H^{\prime}}(X,Y)$ $=e_{H^{\prime}}(X,Y\cup W)-e_{H^{\prime}}(X,W)$
$\geq(e_{H}(X,Y\cup W)-\beta n^{2})-\xi n^{2}$
$\geq\zeta\cdot|X|\cdot|Y|-\beta n^{2}-\xi n^{2}$
$\geq\zeta/2\cdot|X|\cdot|Y|.$
since $\xi,\beta\ll\delta,\zeta$. This shows that $H^{\prime\prime}$ has no $\zeta/2$-sparse cuts, proving (i).
Now we prove (ii). Let $p,p^{\prime}\in V(H^{\prime\prime})$. By (i), $H^{\prime\prime}$ is connected, so it contains a $(p,p^{\prime})$-path. Let $P$ be a shortest $(p,p^{\prime})$-path in $H^{\prime\prime}$. We claim that $P$ has length at most $15/\delta$. Indeed, suppose otherwise, and let $P$ be the path $u_{0}\ldots u_{t}$, where $u_{0}=p$ and $u_{t}=p^{\prime}$ such that $t>15/\delta$. Set $U\coloneqq\{u_{0},u_{3},\ldots,u_{3q}\}$, where $q\coloneqq\lceil 4/\delta\rceil$ (this is well defined because $t>\frac{15}{\delta}\geq 3\cdot(\frac{4}{\delta}+1)\geq 3\cdot\left\lceil\frac{4}{\delta}\right\rceil=3q$). Since $H^{\prime\prime}$ has minimum degree at least $\delta n/4$, the number of edges of $H^{\prime\prime}$ incident to vertices in $U$ is at least $(q+1)\cdot\delta n/4>n$, implying that there is a vertex $v$ in $H^{\prime\prime}$ with at least two neighbours in $U$, say $u_{3i}$ and $u_{3j}$, where $i<j$. Denote by $W$ the walk $u_{0}u_{1}\ldots u_{3i}vu_{3j}\ldots u_{t}$. Note that $W$ is a $(p,p^{\prime})$-walk in $H^{\prime\prime}$. Let $P^{\prime}$ be a $(p,p^{\prime})$-path contained in $W$. Since $W$ is shorter than $P$, it follows that $P^{\prime}$ is also shorter than $P$, a contradiction to the minimality of $P$. This proves (ii), completing the proof of Lemma 7.1. ∎
7.2 Balancing the expanders
As further preparation for the proof of Theorem 1.4, we state the following corollary of Lemma 3.4, which will allow us to balance the expanders using a suitable collection of paths. This is somewhat similar to Lemma 4.1, where we also balance the expanders, but using a collection of vertex-disjoint $K_{t,t}$’s instead of paths.
Lemma 7.2.
Let $\eta,\beta,\xi,\gamma,\zeta,\delta,c\in(0,1)$ and $n\in\mathbb{N}$ satisfy $1/n\ll\eta\ll\beta\ll\xi\ll\gamma\ll\zeta\ll\delta\ll c$. Let $G$ be a $d$-regular graph on $n$ vertices, where $d\geq cn$. Suppose that $\{Z_{1},\ldots,Z_{r}\}$ is a partition of $V(G)$ satisfying properties (E1)(E4) in Lemma 3.2, where $r\leq\left\lceil 1/c\right\rceil$. For $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite, let $\{X_{i},Y_{i}\}$ be a partition of $Z_{i}$ maximising $e_{G}(X_{i},Y_{i})$. Then there are vertex-disjoint paths $P_{1},\ldots,P_{r}$ in $G$ satisfying the following properties, where $Q\coloneqq\bigcup_{i\in[r]}P_{i}$.
  1. (P’1)
    For every $i\in[r]$, we have $2\leq|V(P_{i})|\leq\xi n$.
  2. (P’2)
    For every $i\in[r]$, both leaves of $P_{i}$ are in $Z_{i}$. Moreover, for each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite, one of the leaves of $P_{i}$ is in $X_{i}$ and the other leaf is in $Y_{i}$.
  3. (P’3)
    For each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite, $|X_{i}-V(Q)|=|Y_{i}-V(Q)|$.
Let $H$ be the linear forest guaranteed by Lemma 3.4, satisfying properties (P1)(P4). Let $G^{\prime}$ be the subgraph of $G$ obtained by removing all edges with both endpoints in $X_{i}$ or both endpoints in $Y_{i}$, for each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite.
We will iteratively add some edges from $G^{\prime}$ to $H$ using the following procedure to construct subgraphs $H_{0}\coloneqq H,H_{1},\ldots,H_{r}$ satisfying $|V(H_{i})|\leq\xi n+15i/\delta$ for every $i\in[r]$.
For $i\in[r]$, suppose that $H_{i-1}$ is already defined such that $|V(H_{i-1})|\leq\xi n+15(i-1)/\delta$; we construct $H_{i}$ as follows. If $H_{i-1}$ has two distinct components, say $P$ and $P^{\prime}$, with leaves $p,p^{\prime}$ in $Z_{i}$, say $p\in V(P)\cap Z_{i}$, $p^{\prime}\in V(P^{\prime})\cap Z_{i}$, let $L_{i}$ be a shortest $(p,p^{\prime})$-path in $G^{\prime}[Z_{i}]-(V(H_{i-1})-\{p,p^{\prime}\})$. Applying (ii) of Lemma 7.1 with $G[Z_{i}]$, $G^{\prime}[Z_{i}]$, $V(H_{i-1})-\{p,p^{\prime}\}$, $2\xi$ playing the roles of $H$, $H^{\prime}$, $W$, $\xi$, respectively, we obtain that the path $L_{i}$ has length at most $15/\delta$. (Note that Lemma 7.1 indeed applies because $|V(H_{i-1})|\leq\xi n+15(i-1)/\delta\leq 2\xi n$.) Define $H_{i}\coloneqq H_{i-1}\cup L_{i}$. Note that $|V(H_{i})|\leq|V(H_{i-1})|+15/\delta\leq\xi n+15i/\delta$, as required.
If $G[Z_{i}]$ is $\beta$-almost-bipartite, then, by (P3), one of the two vertices $p,p^{\prime}$ is in $X_{i}$ and the other one is in $Y_{i}$. Therefore, we have the following.
$\text{If $G[Z_{i}]$ is $\beta$-almost-bipartite, then $|V(L_{i})\cap X_{i}|=|V(L_{i})\cap Y_{i}|$}.$ (21)
Notice that for every $i\in[r]$, the number of leaves of $H_{j}$ in $Z_{i}$ is the same as the number of leaves of $H_{j-1}$ in $Z_{i}$, unless $j=i$, in which case it either remains the same, or it decreases from $2$ to 0. Moreover, if $H_{i}$ has two leaves in $Z_{i}$, then they must belong to the same component of $H_{i}$, which is then also a component of $H_{r}$; denote this component by $P_{i}$. For every $i\in[r]$ such that $P_{i}$ is not defined (which means that $H_{r}$ has no components with leaves in $Z_{i}$), define $P_{i}$ to be an edge in $G^{\prime}[Z_{i}]-V(H_{r})$. This is indeed possible because
$|V(H_{r})|\leq\xi n+15r/\delta\leq 2\xi n,$ (22)
which implies that $G^{\prime}[Z_{i}]-V(H_{r})$ contains at least one edge.
Notice that $P_{1},\ldots,P_{r}$ are pairwise vertex-disjoint. We claim that they satisfy (P’1)(P’3) (with $2\xi$ playing the role of $\xi$ for (P’1)). Indeed, by (22) we have $|V(P_{i})|\leq\max\{2,|V(H_{r})|\}\leq 2\xi n$, and since $P_{i}$ contains two leaves, we have $|V(P_{i})|\geq 2$, proving (P’1). Note that (P’2) holds when $P_{i}$ is a component of $H_{r}$ by the discussion in the previous paragraph together with (P3); otherwise, it holds by the choice of $P_{i}$ as an edge in $G^{\prime}[Z_{i}]-V(H_{r})$. For proving (P’3), recall that $Q=P_{1}\cup\ldots\cup P_{r}$ and consider $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite. Also recall that if $L_{i}$ is defined, then its leaves are $p,p^{\prime}$. Then we have,
V(Q)Zi={V(H)Ziif PiHr(V(H)Zi)
V
if PiHr and Li is not defined
(V(H)Zi)
V
(V(Li){p,p})
if PiHr and Li is defined.
[…]
(Here A
B
[…]
denotes the union of the disjoint sets $A$ and $B$.) Hence, (P’3) follows from (P4), (21), and the choice of $P_{i}$ as an edge in $G^{\prime}[Z_{i}]-V(P_{r})$ when $P_{i}\nsubseteq H_{r}$, completing the proof of Lemma 7.2. ∎
7.3 Proof of Theorem 1.7
We are now ready to prove Theorem 1.7.
Let $F$ be a graph with at least one edge, let $n_{0}\in\mathbb{N}$ be such that $1/n_{0}\ll c$, and let $G$ be a $d$-regular graph of order $n\geq n_{0}$ and let $d\geq cn$. Our aim is to show that $G$ has a perfect $\operatorname{TF}$-packing.
By applying Lemma 3.2 to $G$, we obtain positive numbers $\eta,\beta,\gamma,\zeta,\delta$ where
$1/n_{0}\ll\eta\ll\beta\ll\gamma\ll\zeta\ll\delta\ll c$
and a partition $\{Z_{1},\ldots,Z_{r}\}$ of $V(G)$ satisfying properties (E1)(E4) such that $r\leq\lceil 1/c\rceil$.
Let $\xi\in(0,1)$ satisfy
$\beta\ll\xi\ll\gamma.$
For each $i\in[r]$, let $X_{i},Y_{i}$} be a partition of $Z_{i}$ that maximizes $e_{G}(X_{i},Y_{i})$. Let $G^{\prime}$ be the subgraph of $G$ obtained by removing all edges with both endpoints in $X_{i}$ or both in $Y_{i}$, for each $i\in[r]$ such that $G[Z_{i}]$ is $\beta$-almost-bipartite. Let $P_{1},\ldots,P_{r}$ be vertex-disjoint paths satisfying (P’1)(P’3), guaranteed by Lemma 7.2, and let $Q\coloneqq P_{1}\cup\ldots\cup P_{r}$.
7.3.1 Constructing a balanced pair of $F$-subdivisions in each expander
For every $i\in[r]$, we will construct two vertex-disjoint subdivisions of $F$ in $G^{\prime}[Z_{i}]-V(Q)$ as follows.
Claim 7.3.
For every $i\in[r]$, there exist two vertex-disjoint $F$-subdivisions $F_{i},F^{\prime}_{i}$ in $G^{\prime}[Z_{i}]-V(Q)$ such that $|V(F_{i})|,|V(F_{i}^{\prime})|\leq e(F)\cdot 15/\delta$. Moreover, if $G[Z_{i}]$ is $\beta$-almost-bipartite, then
$|X_{i}\cap(V(F_{i}\cup F^{\prime}_{i}))|=|Y_{i}\cap(V(F_{i}\cup F^{\prime}_{i}))|.$ (23)
Fix $i\in[r]$. Let $S,S^{\prime}\subseteq Z_{i}-V(Q)$ be disjoint sets of size $|V(F)|$, where $S\subseteq X_{i}$ and $S^{\prime}\subseteq Y_{i}$ if $G[Z_{i}]$ is $\beta$-almost-bipartite, and let $\phi:S\to V(F)$ and $\phi^{\prime}:S^{\prime}\to V(F)$ be two injections chosen arbitrarily. Let $T$ be the set of unordered pairs $\{s,s^{\prime}\}$ with $s,s^{\prime}\in S$ such that $\phi(s)\phi(s^{\prime})$ is an edge in $F$, and similarly, let $T^{\prime}$ be the set of unordered pairs $\{s,s^{\prime}\}$ with $s,s^{\prime}\in S^{\prime}$ such that $\phi^{\prime}(s)\phi^{\prime}(s^{\prime})$ is an edge in $F$.
Let $\mathcal{P}$ be a maximal collection of pairwise internally vertex-disjoint paths in $G^{\prime}[Z_{i}]-V(Q)$, with internal vertices in $Z_{i}-(V(Q)\cup S\cup S^{\prime})$, whose length is at most $15/\delta$ and such that each of the paths in $\mathcal{P}$ is an $(s,s^{\prime})$-path for a distinct pair $\{s,s^{\prime}\}\in T\cup T^{\prime}$. We claim that $|\mathcal{P}|=2e(F)$. Indeed, suppose otherwise. Then there is a pair $(s,s^{\prime})\in T\cup T^{\prime}$ for which there is no $(s,s^{\prime})$-path in $\mathcal{P}$. Fix such a pair $(s,s^{\prime})$, and let $W\coloneqq V(\mathcal{P})\cup V(Q)\cup S\cup S^{\prime}$. Then
$|W|\leq 2e(F)\cdot 15/\delta+\xi n+2|V(F)|\leq 2\xi n.$
Thus, by applying (ii) of Lemma 7.1 (with $2\xi$ playing the role of $\xi$), we obtain an $(s,s^{\prime})$-path $P$ in $G^{\prime}[Z_{i}]-W$ of length at most $15/\delta$. But then $\mathcal{P}\cup\{P\}$ contradicts the maximality of $\mathcal{P}$. This shows that $|\mathcal{P}|=2e(F)$, as claimed.
Pick a collection $\mathcal{P}$ as guaranteed by the previous paragraph. Take $F_{i}$ to be the subgraph consisting of the paths in $\mathcal{P}$ with ends in $S$, and take $F_{i}^{\prime}$ to be the subgraph consisting of the paths in $\mathcal{P}$ with ends in $S^{\prime}$. We claim that $F_{i}$ and $F_{i}^{\prime}$ satisfy the requirements of Claim 7.3. Indeed, first note that $F_{i}$ is an $F$-subdivision in $G^{\prime}[Z_{i}]-V(Q)$ whose branch vertices are in $S$, and similarly $F_{i}^{\prime}$ is an $F$-subdivision in $G^{\prime}[Z_{i}]-V(Q)$ whose branch vertices are in $S^{\prime}$. Second, notice that $F_{i}$ and $F_{i}^{\prime}$ are vertex-disjoint, since the paths in $\mathcal{P}$ are internally vertex-disjoint whose internal vertices are in $Z_{i}-(V(Q)\cup S\cup S^{\prime})$. Third, we have $|V(F_{i})|,|V(F_{i}^{\prime})|\leq e(F)\cdot 15/\delta$, since every path in $\mathcal{P}$ has length at most $15/\delta$ by our choice of $\mathcal{P}$. Finally, if $G[Z_{i}]$ is $\beta$-almost-bipartite, then for every path $P\in\mathcal{P}$ with ends in $S$, we have
$|V(P^{\circ})\cap X_{i}|=|V(P^{\circ})\cap Y_{i}|-1,$
where $P^{\circ}$ denotes the interior of the path $P$. Hence
[…]
$=-e(F)+|V(F)|,$
and analogously,
$|V(F_{i}^{\prime})\cap Y_{i}|-|V(F_{i}^{\prime})\cap X_{i}|=-e(F)+|V(F)|.$
Altogether, since $F_{i}$ and $F_{i}^{\prime}$ are vertex-disjoint,
$|V(F_{i}\cup F_{i}^{\prime})\cap X_{i}|-|V(F_{i}\cup F_{i}^{\prime})\cap Y_{i}|$
$\qquad=|V(F_{i})\cap X_{i}|-|V(F_{i})\cap Y_{i}|+|V(F_{i}^{\prime})\cap X_{i}|-|V(F_{i}^{\prime})\cap Y_{i}|=0.$
This proves (23), completing the proof of Claim 7.3. ∎
7.3.2 Absorbing the paths $P_{1},\ldots,P_{r}$ and all the uncovered vertices into the subdivisions $F_{i}$
For every $i\in[r]$, let $F_{i},F^{\prime}_{i}$ be the two vertex-disjoint subdivisions of $F$ in $G^{\prime}[Z_{i}]-V(Q)$, as guaranteed by Claim 7.3. Denote the leaves of $P_{i}$ by $u_{i}$ and $v_{i}$, where $u_{i}\in X_{i}$ and $v_{i}\in Y_{i}$ if $G[Z_{i}]$ is $\beta$-almost-bipartite. Let $x_{i}y_{i}$ be an arbitrary edge of $F_{i}$, where $x_{i}\in X_{i}$ and $y_{i}\in Y_{i}$ if $G[Z_{i}]$ is $\beta$-almost-bipartite. Let $Q_{i}$ be a shortest $(y_{i},u_{i})$-path in $G^{\prime}[Z_{i}]$ whose interior avoids $V(Q)\cup V(F_{i})\cup V(F_{i}^{\prime})$. By (P’1),
$|V(Q)|+|V(F_{i})|+|V(F_{i}^{\prime})|\leq r\xi n+2|V(F)|\leq\lceil 1/c\rceil\xi n+2|V(F)|\leq\sqrt{\xi}n.$ (24)
Hence, it follows that $Q_{i}$ exists and is of length at most $15/\delta$ by applying (ii) of Lemma 7.1 (with $\sqrt{\xi}$ playing the role of $\xi$). Let $H\coloneqq G[Z_{i}]$, and let $H^{\prime}\coloneqq G^{\prime}[Z_{i}]$. Then, by Lemma 7.1, $H^{\prime}$ has no $\zeta/2$-sparse cuts. Now let
$W_{i}\coloneqq Z_{i}\cap\Big(\big(V(Q)\cup V(F_{i})\cup V(F_{i}^{\prime})\cup V(Q_{i})\big)-\{x_{i},v_{i}\}\Big).$
By (24) and the fact that $|V(Q_{i})|\leq\frac{15}{\delta}+1$, we have $|W_{i}|\leq\frac{15}{\delta}+1+\sqrt{\xi}n\leq 2\sqrt{\xi}n.$
Note that $H^{\prime}$ has maximum degree at most $d$ since it is a subgraph of $G$. By (E1), and the fact that $|V(H^{\prime})|=|Z_{i}|\geq\delta n$ (by (E2)), we have
$2e(H^{\prime})\geq d\cdot|V(H^{\prime})|-\eta n^{2}-\beta n^{2}\geq d\cdot|V(H^{\prime})|-\frac{2\beta}{\delta^{2}}\cdot|V(H^{\prime})|^{2},$
showing that $H^{\prime}$ has average degree at least $d-(2\beta/\delta^{2})|V(H^{\prime})|$. Moreover, $H^{\prime}$ is either bipartite or $\gamma$-far-from-bipartite by definition. Therefore, using $|W_{i}|\leq 2\sqrt{\xi}n$, we can apply Lemma 3.3 with $H^{\prime}$, $W_{i}$, $2\beta/\delta^{2}$, $2\xi$, $\zeta/2$ playing the roles of $G$, $W$, $\eta$, $\xi$, $\zeta$, respectively, to find a Hamilton path $Q_{i}^{\prime}$ in $H^{\prime}-W_{i}=G^{\prime}[Z_{i}]-W_{i}$ with ends $x_{i}$ and $v_{i}$. Note that Lemma 3.3 indeed applies because when $G[Z_{i}]$ is $\beta$-almost-bipartite (so $H^{\prime}$ is bipartite), (P’3) and (23) together imply that $|X_{i}-W_{i}|=|Y_{i}-W_{i}|$.
Now replace the edge $x_{i}y_{i}$ in $F_{i}$ with the path $x_{i}Q_{i}^{\prime}v_{i}P_{i}u_{i}Q_{i}y_{i}$ to obtain a subgraph $F^{\prime\prime}_{i}$ which is a subdivision of $F$. Note that $V(F^{\prime\prime}_{i})\cup V(F^{\prime}_{i})=(Z_{i}-V(Q))\cup V(P_{i})$. Hence, $\{F^{\prime}_{i},F^{\prime\prime}_{i}\mid i\in[r]\}$ is a collection of $F$-subdivisions that are pairwise vertex-disjoint covering all vertices of $G$, completing the proof of Theorem 1.7. ∎
8 Concluding remarks
Recall that our main result (Theorem 1.4) states that for every bipartite graph $H$, every dense regular graph $G$ contains an $H$-packing covering all but $O(1)$ vertices. As noted in the introduction, this was already known for unbalanced bipartite graphs $H$ by Kühn and Osthus [31] (see Theorem 1.3). In fact, they proved a slightly stronger statement in this case: the host graph $G$ need only be almost regular rather than regular. More precisely, for every bipartite $H$ and every $c>0$, there exist $\varepsilon>0$ and $C$ such that any $n$-vertex graph $G$ in which every vertex has degree between $(c-\varepsilon)n$ and $(c+\varepsilon)n$ contains an $H$-packing covering all but at most $C$ vertices. This fails, however, for balanced bipartite $H$ (for example, if $G$ is a slightly unbalanced bipartite graph) so our result cannot, in general, be extended to the almost regular setting.
Keevash proved an analogous result for hypergraphs (see Theorem 5.14 in [15]). [^2] Keevash’s theorem is stated and proved for $3$-uniform hypergraphs, but the analogous statement for higher uniformities can be obtained by the same methods. Specifically, he showed that for every $3$-partite 33-graph $H$ whose parts are not all of equal size, and for every $0<c_{1},c_{2}<1$, there exist $\varepsilon>0$ and $C>0$ such that if $G$ is a $3$-graph on $n$ vertices in which every vertex has degree between $(1-\varepsilon)c_{1}n^{2}$ and $(1+\varepsilon)c_{1}n^{2}$, and every pair of vertices has degree more than $c_{2}n$, then $G$ contains an $H$-packing covering all but at most $C$ vertices. Here, the degree of a pair $\{u,v\}$ of vertices in $G$ is the number of vertices $w\in V(G)-\{u,v\}$ such that $\{u,v,w\}\in E(G)$.
It is plausible that, analogous to Theorem 1.4, the conclusion holds for all tripartite $3$-graphs $H$ when $G$ is assumed to be regular.
Question 8.1.
Is it true that for every tripartite $3$-graph $H$ and every $0<c<1$, there exist $\varepsilon>0$ and $C>0$ such that the following holds? If $G$ is a $3$-graph on $n$ vertices in which all vertices have the same degree, which is at least $cn^{2}$, and every pair of vertices has degree at least $cn$, then $G$ contains an $H$-packing covering all but at most $C$ vertices.
The problem of partitioning the vertex set of edge-coloured complete graphs into a small number of monochromatic subgraphs has a very rich history; see [12] for a recent survey. An early example of a problem of this kind is Lehel’s conjecture. An interesting problem in this direction, suggested by Matija Bucić, is to consider a variant of Theorem 1.4 where the edges of $G$ are coloured and we seek packings using monochromatic $H$-copies (where different copies of $H$ may receive different colours).
Question 8.2.
Is it true that for every bipartite graph $H$, integer $r\geq 2$, and $0<c<1$, there exists $C>0$ such that the following holds? If $G$ is an $n$-vertex $d$-regular graph with $d\geq cn$, whose edges are coloured with $r$ colours, then $G$ contains a collection of vertex-disjoint monochromatic copies of $H$ covering all but at most $C$ vertices.
The problem of finding a large monochromatic $H$-packing in a graph with a given minimum degree was studied by Balogh, Freschi, and Treglown [2].
Appendix A Perfect fractional matching implies perfect $2$-matching
Suppose that $G$ is a graph with a perfect fractional matching. Let $w:E(G)\mapsto[0,1]$ be a perfect fractional matching in $G$ that minimises the number of edges whose weight is neither 0 nor $1$, and denote by $G^{\prime}$ the subgraph of $G$ whose edges have non-zero weight in $w$. We claim that $G^{\prime}$ is a $2$-matching; that is, every connected component of $G^{\prime}$ is either a single edge or an odd cycle. Notice that $G^{\prime}$ is spanning, because $w$ is a perfect fractional matching. Thus, if $G^{\prime}$ is a $2$-matching, this would show that $G^{\prime}$ is a perfect $2$-matching in $G$ as required.
Therefore, it remains to show that $G^{\prime}$ is a $2$-matching. Suppose for a contradiction that it is not, and let $F$ be a component of $G^{\prime}$ which is neither an edge nor an odd cycle.
First, notice that $F$ has minimum degree at least $2$. Indeed, otherwise, let $v$ be a vertex with degree $1$ in $F$, let $u$ be its unique neighbour, and let $x$ be a neighbour of $u$ which is not $v$. Then $w(vu),w(ux)>0$, implying that $w(vu)<1$, contradicting the assumption that $w$ is a perfect fractional matching.
Next, we claim that $F$ does not have even cycles. Indeed, suppose that $C=(v_{1}\ldots v_{2s})$ is an even cycle in $F$. For $x>0$, let $w_{x}:E(G)\mapsto[0,1]$ be the weighting of the edges of $G$ defined as follows (with the addition of indices taken modulo $2s$).
$w_{x}=\left\{\begin{array}[]{ll}w(e)&e\notin E(C)\\ w(e)+x&e=v_{2i-1}v_{2i}\text{ for $i\in[s]$}\\ w(e)-x&e=v_{2i}v_{2i+1}\text{ for $i\in[s]$}.\end{array}\right$.
Note that the edges $v_{i}v_{i+1}$, with $i\in[2s]$, have weight strictly between 0 and $1$, and so there exists $x>0$ such that all weights in $w_{x}$ are in $[0,1]$. Let $x^{\prime}$ be the maximum $x$ with this property. Then at least one of the edges in $C$ has weight either 0 or $1$ in $w_{x^{\prime}}$, showing that $w_{x^{\prime}}$ is a perfect fractional matching in $G$ with fewer edges whose weight is neither 0 nor $1$ compared to $w$, a contradiction to the choice of $w$. This shows that $F$ does not have even cycles, as desired. Moreover, this shows that $F$ cannot be a cycle since we assumed that $F$ is not an odd cycle.
Since $F$ has minimum degree at least $2$ and is not a cycle, it contains two distinct cycles $C_{1}$ and $C_{2}$. Because $F$ has no even cycles, these cycles are odd and share at most one vertex (otherwise their union contains an even cycle). Let $P$ be a shortest path with one end in $C_{1}$ and the other end in $C_{2}$. (Note that $P$ is a singleton if the cycles share a vertex.) Write $C_{1}=(v_{1}\ldots v_{2s+1})$, $C_{2}=(u_{1}\ldots u_{2t+1})$, and suppose that $P$ is a path from $v_{1}$ to $u_{1}$. Now, for $x>0$, let $w_{x}:E(G)\mapsto[0,1]$ be the weighting of the edges of $G$ obtained from $w$ by making the following modifications.
  • Decrease the weight of edges $v_{2i-1}v_{2i}$, with $i\in[s+1]$, by $x$ (where addition in the indices is taken modulo $2s+1$).
  • Increase the weight of edges $v_{2i}v_{2i+1}$, with $i\in[s]$, by $x$.
  • Increase the weight of the first, third, fifth, etc. edges of $P$ by $2x$ (starting from $v_{1}$).
  • Decrease the weight of the second, fourth, sixth, etc. edges of $P$ by $2x$ (starting from $v_{1}$).
  • If $P$ has odd length, decrease the weight of edges $u_{2i-1}u_{2i}$ with $i\in[t+1]$ by $x$ (where addition of indices is taken modulo $2t+1$). Otherwise, increase the weight of these edges by $x$.
  • If $P$ has odd length, increase the weight of edges $u_{2i}u_{2i+1}$, with $i\in[t]$, by $x$. Otherwise, decrease the weight of these edges by $x$.
As before, there exists $x>0$ such that all weights in $w_{x}$ are in $[0,1]$, and let $x^{\prime}$ be the maximum $x$ with this property. Then $w_{x^{\prime}}$ is a perfect fractional matching of $G$ with fewer edges whose weight is neither 0 nor $1$ compared to $w$, a contradiction.
In conclusion, the assumption that $F$ is a component of $G^{\prime}$ that is neither an odd cycle nor an edge leads to a contradiction, so $G^{\prime}$ is a $2$-matching (and thus $G^{\prime}$ is a perfect $2$-matching in $G$), as required. ∎
Appendix B Hamiltonicity of clusters
In this section, we present the proof of Lemma 3.3. Our proof closely follows that of Lemma 4 in [11], but we include it here for completeness and to address differences in the lemma’s statement.
The proof is based on known results concerning the Hamiltonicity of robust out-expanders, a concept introduced by Kühn, Osthus and Treglown [34]. Before stating the relevant result, we first introduce some necessary definitions.
Given a digraph $G$ on $n$ vertices, a set of vertices $S$ and a parameter $\nu\in(0,1)$, the robust $\nu$-out-neighbourhood of $S$ in $G$, denoted $\mathrm{RN}^{+}_{\nu,G}(S)$, is the set of vertices in $G$ that have at least $\nu n$ in-neighbours in $S$; we omit the subscript $G$ when it is clear from the context. Given $0<\nu\leq\tau<1$, we say that $G$ is a robust $(\nu,\tau)$-out-expander if $|\mathrm{RN}^{+}_{\nu}(S)|\geq|S|+\nu n$ for every set of vertices $S$ with $\tau n\leq|S|\leq(1-\tau)n$. We shall also use the following undirected version of a robust out-neighbourhood. In a graph $G$ on $n$ vertices, the robust $\nu$-neighbourhood of a set of vertices $S$, denoted $\mathrm{RN}_{\nu,G}(S)$, is the set of vertices in $G$ with at least $\nu n$ neighbours in $S$. As before, we may omit the subscript $G$ when it is clear from the context. We say that $G$ is a robust $(\nu,\tau)$-expander if $|\mathrm{RN}_{\nu,G}(S)|\geq|S|+\nu n$ for every set of vertices $S$ with $\tau n\leq|S|\leq(1-\tau)n$.
We will need the following theorem from [34]. The minimum semi-degree of a digraph $G$, denoted $\delta^{0}(G)$, is defined as $\delta^{0}(G)=\min\{\delta^{+}(G),\delta^{-}(G)\}$, where $\delta^{+}(G),\delta^{-}(G)$ are the minimum out-degree and in-degree of $G$, respectively.
Theorem B.1 (Kühn, Osthus and Treglown [34]).
Let $n\in\mathbb{N}$ and let $\nu,\tau,\gamma$ be reals such that $1/n\ll\nu\leq\tau\ll\gamma<1$. Let $G$ be a digraph on $n$ vertices with $\delta^{0}(G)\geq\gamma n$ which is a robust $(\nu,\tau)$-out-expander. Then $G$ contains a Hamilton cycle.
We shall use the following corollary of Theorem B.1.
Corollary B.2.
Let $\nu,\tau,\gamma$ be reals and let $n$ be an integer such that $1/n\ll\nu\leq\tau\ll\gamma<1$. Let $G$ be a digraph on $n$ vertices with $\delta^{0}(G)\geq\gamma n$ which is a robust $(\nu,\tau)$-out-expander. Then for every choice of distinct vertices $x$ and $y$, there is a Hamilton path in $G$ with ends $x$ and $y$.
Given vertices $x$ and $y$, form $G^{\prime}$ by adding the edge $xy$ to $G$, removing the edge $yx$ (if it exists), and removing all the edges directed toward $y$ or away from $x$. Next, form $G^{\prime\prime}$ by contracting the edge $xy$. It is easy to check that $G^{\prime\prime}$ is a robust $(\nu/2,2\tau)$-out-expander. Thus, by Theorem B.1, it contains a Hamilton cycle. This cycle corresponds to a Hamilton cycle in $G^{\prime}$ which contains the edge $xy$, which in turn corresponds to a Hamilton path in $G$ with ends $x$ and $y$. ∎
Let $G$ be a graph with $n$ vertices, maximum degree at most $d$, and average degree at least $d-\eta n$. Suppose $G$ has no $\zeta$-sparse cuts and is either $\gamma$-far-from-bipartite or bipartite with a bipartition $\{X,Y\}$. Let $W\subseteq V(G)$ be a set of size at most $\xi n$, which satisfies |XW|=|YW||X-W|=|Y-W| if $G$ is bipartite. Define H=GWH=G-W, $V=V(H)$, and, if $G$ is bipartite, then also define $X^{\prime}=X-W$ and $Y^{\prime}=Y-W$. Then $H$ has maximum degree at most $d$, average degree at least $d-2\xi n$, and no $\zeta/2$-sparse cuts.
The following claim will allow us to use Corollary B.2 above.
Claim B.3.
Let $S\subseteq V$ be a set satisfying $\xi^{1/7}n\leq|S|\leq(1-\xi^{1/7})n$ if $G$ is $\gamma$-far-from-bipartite, or $\xi^{1/7}n\leq|S|\leq(1/2-\xi^{1/7})n$ if $G$ is bipartite. Then $|\mathrm{RN}_{\xi,H}(S)|\geq|S|+\xi n$.
Define $S_{1}=S-\mathrm{RN}_{\xi}(S)$, $S_{2}=S\cap\mathrm{RN}_{\xi}(S)$, $T_{1}=\mathrm{RN}_{\xi}(S)-S$, $T_{2}=V-(S\cup T_{1})$. Assume, towards a contradiction, that $|\mathrm{RN}_{\xi}(S)|<|S|+\xi n$, which implies that $|T_{1}|<|S_{1}|+\xi n$.
Given sets $A,B\subseteq V$, denote by $e(A,B)$ the number of ordered pairs $ab$ such that $ab$ is an edge of $H$ and $a\in A,b\in B$. Then
$e(S_{1},V-T_{1})\leq e(S_{1},S)+e(S_{1},T_{2})\leq(|S_{1}|+|T_{2}|)\cdot\xi n\leq\xi n^{2},$ (25)
using that the vertices in $S_{1}\cup T_{2}$ are not in $\mathrm{RN}_{\xi}(S)$. By the maximum and average degree assumption on $H$, we have
$e(S_{1},T_{1})=\sum_{v\in S_{1}}d_{H}(v)-e(S_{1},V-T_{1})\geq d|S_{1}|-2\xi n^{2}-\xi n^{2}=d|S_{1}|-3\xi n^{2}.$
As $|T_{1}|<|S_{1}|+\xi n$, we obtain the following bound on $e(T_{1},V-S_{1})$.
$e(T_{1},V-S_{1})\leq d|T_{1}|-e(S_{1},T_{1})\leq d(|T_{1}|-|S_{1}|)+3\xi n^{2}\leq 4\xi n^{2}.$ (26)
Consider the quantity $e(S_{1}\cup T_{1},V-(S_{1}\cup T_{1}))$. By (25) and (26), this quantity is at most $5\xi n^{2}$, and since $H$ has no $\zeta/2$-sparse cuts, it is at least $(\zeta/2)\cdot|S_{1}\cup T_{1}|\cdot|V-(S_{1}\cup T_{1})|$. As $\xi\ll\zeta$, we find that either $|S_{1}\cup T_{1}|\leq\xi^{1/3}n$ or $|V-(S_{1}\cup T_{1})|\leq\xi^{1/3}n$.
First, suppose that $|S_{1}\cup T_{1}|\leq\xi^{1/3}n$. Then
$e(S_{2},V-S_{2})\leq e(S_{1}\cup T_{1},V)+e(S_{2},T_{2})\leq\xi^{1/3}n^{2}+|T_{2}|\xi n\leq 2\xi^{1/3}n^{2},$
using that $|S_{1}\cup T_{1}|\leq\xi^{1/3}n$ and $T_{2}\cap\mathrm{RN}_{\xi}(S_{2})=\emptyset$. However, since $H$ does not have $\zeta/2$-sparse cuts and given the assumption of the size of $S$, we have
$e(S_{2},V-S_{2})\geq(\zeta/2)\cdot|S_{2}|\cdot|V-S_{2}|\geq\frac{\zeta}{2}\cdot(\xi^{1/7}n-|S_{1}|)\cdot\frac{\xi^{1/7}n}{2}\geq\frac{\zeta}{2}\cdot\frac{\xi^{2/7}n^{2}}{4}>2\xi^{1/3}n^{2},$
a contradiction.
Next, suppose that $|V-(S_{1}\cup T_{1})|\leq\xi^{1/3}n$. If $G$ is bipartite, then using that $|T_{1}|<|S_{1}|+\xi n$ and $|S|\leq(1/2-\xi^{1/7})n$, we have
$|V|=|V-(S_{1}\cup T_{1})|+|S_{1}\cup T_{1}|$ $<\xi^{1/3}n+|S_{1}|+(|S_{1}|+\xi n)$
$\leq\xi^{1/3}n+n-2\xi^{1/7}n+\xi n<n-\xi n\leq n-|W|,$
a contradiction. Hence, we may assume that $G$ is $\gamma$-far-from-bipartite. Note that $G$ can be made bipartite by removing edges incident to $W\cup S_{2}\cup T_{2}$ or within $S_{1}$ or $T_{1}$. However, there are at most $(\xi+\xi^{1/3})n^{2}$ edges of the former type (since $|W|\leq\xi n$ and $|V-(S_{1}\cup T_{1})|=|S_{2}\cup T_{2}|\leq\xi^{1/3}n$) and at most $5\xi n^{2}$ edges of the latter type (by (25) and (26)), so there are fewer than $\gamma n^{2}$ edges in total (using $\xi\ll\gamma$). This contradicts the assumption that $G$ is $\gamma$-far-from-bipartite, thus completing the proof of the claim. ∎
It follows from Claim B.3 that $H$ is a robust $(\xi,2\xi^{1/7})$-expander. Let $x,y\in V$, where $x\in X$ and $y\in Y$ if $G$ is bipartite. Our goal is to show that $H$ contains a Hamilton path with ends $x$ and $y$. First, consider the case where $G$ is $\gamma$-far-from-bipartite. Form a digraph $D$ by replacing each edge $uv$ of $H$ with the two directed edges $uv$ and $vu$. It follows from Claim B.3 that $D$ is a robust $(\xi,2\xi^{1/7})$-out-expander. Then, Corollary B.2 implies the existence of a Hamilton path in $D$ with ends $x$ and $y$, which corresponds to a Hamilton path in $H$ with the same ends.
Now, suppose that $G$ is bipartite. Note that in this case $H$ is a balanced bipartite graph (by our choice of the set $W$). We first show that $H$ has a perfect matching by verifying Hall’s condition i.e., we will show that for every $S\subseteq X^{\prime}$, we have $|N_{H}(S)|\geq|S|$. Notice that $H$ has minimum degree at least $\zeta(n-1)-\xi n\geq\zeta n/2$, by the assumption that $G$ has no $\zeta$-sparse cuts and by $|W|\leq\xi n$. Thus, if $|S|<\zeta n/2$ (and $S\neq\emptyset$) then $|N_{H}(S)|\geq\zeta n/2\geq|S|$. Similarly, if $|S|>|V|/2-\zeta n/2$ then $|N_{H}(S)|=|V|/2\geq|S|$. Hence, we may assume that $\zeta n/2\leq|S|\leq|V|/2-\zeta n/2$, which implies that $\xi^{1/7}n\leq|S|\leq(1/2-\xi^{1/7})n$, but then Claim B.3 implies that $|N_{H}(S)|\geq|S|$, showing that $H$ has a perfect matching.
Let $\{a_{1}b_{1},\ldots,a_{t}b_{t}\}$ be a perfect matching in $H$, where $t=|X^{\prime}|$ and $a_{i}\in X^{\prime},b_{i}\in Y^{\prime}$ for $i\in[t]$. We assume for convenience that $a_{i}b_{i}$ is not the edge $xy$ (if the edge $xy$ exists) for $i\in[t]$ – this is possible because removal of the edge $xy$ from $H$ does not affect the arguments above. Without loss of generality, $a_{1}=x$ and $b_{t}=y$. Form an auxiliary directed graph $D$ with vertex set $\{v_{1},\ldots,v_{t}\}$ where $v_{i}v_{j}$ is a directed edge whenever $b_{i}a_{j}$ is an edge of $H$. It follows from Claim B.3 that $D$ is a robust $(\xi,3\xi^{1/7})$-out-expander. Then, by Corollary B.2, there is a Hamilton path in $D$ starting at $v_{1}$ and ending at $v_{t}$. Without loss of generality, suppose that this path is $v_{1},\ldots,v_{t}$. This path corresponds to the Hamilton path $x=a_{1},b_{1},\ldots,a_{t},b_{t}=y$ in $H$, as desired. ∎